### v2

# SketchGuard: Scaling Byzantine-Robust Decentralized Federated Learning via Sketch-Based Screening

Abstract.
Decentralized Federated Learning (DFL) enables privacy-preserving collaborative training without centralized servers, but remains vulnerable to Byzantine attacks where malicious clients submit corrupted model updates. Existing Byzantine-robust DFL defenses rely on similarity-based neighbor screening that requires every client to exchange and compare complete high-dimensional model vectors with all neighbors in each training round, creating prohibitive communication and computational costs that prevent deployment at web scale. We propose SketchGuard, a general framework that decouples Byzantine filtering from model aggregation through sketch-based neighbor screening. SketchGuard compresses $d$-dimensional models to $k$-dimensional sketches ($k\ll d$) using Count Sketch for similarity comparisons, then selectively fetches full models only from accepted neighbors, reducing per-round communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$, where $|\mathcal{N}_{i}|$ is the neighbor count and $|\mathcal{S}_{i}|\leq|\mathcal{N}_{i}|$ is the accepted neighbor count. We establish rigorous convergence guarantees in both strongly convex and non-convex settings, proving that Count Sketch compression preserves Byzantine resilience with controlled degradation bounds where approximation errors introduce only a $(1+O(\epsilon))$ factor in the effective threshold parameter. Comprehensive experiments across multiple datasets, network topologies, and attack scenarios demonstrate that SketchGuard maintains identical robustness to state-of-the-art methods, with mean test error rate deviation of only up to 0.35 percentage points, while reducing computation time by up to 82% and communication overhead by 50-70% depending on filtering effectiveness, with benefits scaling multiplicatively with model dimensionality and network connectivity. These results establish the viability of sketch-based compression as a fundamental enabler of robust DFL at web scale.
Decentralized Learning, Byzantine Robustness, Dimensionality Reduction, Scalable Machine Learning
copyright: rightsretainedjournalyear: 2026ccs: Security and privacy Systems security
1. Introduction
Federated Learning (FL) enables collaborative training of AI models over distributed data while preserving privacy by keeping raw data local (McMahan et al., 2017). However, the canonical, server-assisted architecture of FL centralizes aggregation of model parameters, creating a single point of failure, a communication bottleneck, and trust issues (Fang et al., 2024). These drawbacks have catalyzed Decentralized Federated Learning (DFL), where clients exchange model updates in a peer-to-peer manner over a dynamic communication graph, thereby improving scalability and resilience (Beltrán et al., 2023).
A central challenge in DFL is byzantine robustness: the ability to withstand malicious clients that submit arbitrary or carefully crafted updates to poison training, induce consensus drift, or trigger targeted failures (Blanchard et al., 2017; Baruch et al., 2019). Unlike centralized FL where robust rules such as Krum (Blanchard et al., 2017), coordinatewise Median (Yin et al., 2018), or Trimmed-Mean (Wang et al., 2025) are applied once at a centralized server, DFL requires every client to aggregate its neighbors’ updates under local, graph-limited views, often with non-IID data and time-varying connectivity. To address this added complexity, most DFL defenses adopt local-consistency filters, where clients accept a neighbor’s update only if it is sufficiently similar to their own state, and then average over the accepted subset (Guo et al., 2021; He et al., 2022; Fang et al., 2024; Cajaraville-Aboy et al., 2024; El-Mhamdi et al., 2021). These mechanisms provide convergence and robustness guarantees in both strongly convex and non-convex model training settings, yet suffer from a fundamental scalability bottleneck: clients must exchange and compare complete, high-dimensional model vectors with all neighbors in every round. For emerging web-scale applications like decentralized training of frontier models with billions of parameters across thousands of distributed participants (Borzunov et al., 2023; Long, 2024), this creates prohibitive communication and computation costs that prevent practical implementation of these systems at scale.
Sketch-based compression offers tools for communication efficient learning. Count Sketch, for instance, compresses a vector of $d$ dimensions into a summary of $k$ dimensions using simple hash and sign functions, where $k\ll d$. Sketches are linear and allow approximate preservation of coordinates, thereby supporting fast similarity estimation with formal guarantees (Charikar et al., 2002; Rothchild et al., 2020; Gattani et al., 2024). While sketching has proven effective for bandwidth reduction in FL, extending it to Byzantine-robust, fully decentralized aggregation, where compressed representations must support secure neighbor screening, remains underexplored.
In this paper, we propose SketchGuard, a general framework for Byzantine-robust DFL that decouples filtering from aggregation through sketch-based neighbor screening. Our key insight is that similarity-based Byzantine filtering can operate on compressed representations, while the final aggregation requires full precision models only for accepted neighbors. SketchGuard is applicable to any similarity-based Byzantine defense that relies on Euclidean distance measures (e.g., (Fang et al., 2024; He et al., 2022; Sun et al., 2024; Pillutla et al., 2022)), but for our theoretical analysis and empirical evaluation, we instantiate it with state-of-the-art BALANCE aggregation (Fang et al., 2024), which provides the strongest theoretical guarantees among existing methods. Our main contributions can be summarized as follows:
  • We provide rigorous analysis showing that Count Sketch compression maintains Byzantine resilience with controlled degradation bounds.
  • We establish convergence rates for SketchGuard in both strongly convex and non-convex settings.
  • Through comprehensive experiments across multiple datasets, network topologies, and attack scenarios, we demonstrate that our approach achieves identical robustness to state-of-the-art methods while reducing communication overhead by 50-70% and computation time by up to 82%.
2. Preliminaries and Related Work
This section provides technical background on DFL protocols, Byzantine attack models, and compression techniques, and discusses related work in each area.
2.1. DFL Problem Formulation and Protocol
Consider $n$ clients connected by an undirected graph $G=(V,E)$, where each client $i\in V$ possesses a private dataset $\mathcal{D}_{i}$ and maintains a local model $\mathbf{w}_{i}\in\mathbb{R}^{d}$. The collective objective is to minimize the average empirical loss:
(1) $\min_{\mathbf{w}\in\mathbb{R}^{d}}F(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{w}),$
where $f_{i}(\mathbf{w})=\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{i}}[\ell(\mathbf{w};\mathbf{x},y)]$ is the expected loss over client $i$’s data distribution.
The DFL protocol alternates between two phases. First, clients perform local updates:
(2) $\mathbf{w}_{i}^{t+1/2}=\mathbf{w}_{i}^{t}-\eta\nabla f_{i}(\mathbf{w}_{i}^{t}).$
Then, each client aggregates neighbor models according to:
(3) $\mathbf{w}_{i}^{t+1}=\alpha\mathbf{w}_{i}^{t+1/2}+(1-\alpha)\cdot\mathrm{AGG}_{i}\left(\{\mathbf{w}_{j}^{t+1/2}:j\in\mathcal{N}_{i}\}\right),$
where $\mathcal{N}_{i}$ denotes client $i$’s neighbors and $\mathrm{AGG}_{i}$ is a local aggregation function.
2.2. Byzantine Attack Models
We consider $f$-Byzantine adversaries that control up to $f$ clients per neighborhood. Byzantine clients can deviate arbitrarily from the protocol, presenting a spectrum of threats ranging from simple disruption to sophisticated coordinated manipulation. These adversaries may employ model poisoning by sending crafted parameters $\hat{\mathbf{w}}_{j}$ designed to maximize deviation from honest updates while evading detection by robust aggregation rules. Representative attacks include random noise injection (Blanchard et al., 2017), data label flipping (Baruch et al., 2019), and gradient-based perturbation optimization (Muñoz-González et al., 2017). More sophisticated adversaries operate adaptively, observing honest client behavior and adjusting their strategy accordingly. Such attacks (Fang et al., 2020; Shejwalkar and Houmansadr, 2021) solve optimization problems to craft maximally harmful updates that remain within detection thresholds of similarity-based defenses. The most challenging scenarios involve collusive attacks where multiple Byzantine clients coordinate their malicious behavior (Xie et al., 2020; Muñoz-González et al., 2017), potentially overwhelming local robust aggregation mechanisms that assume independent adversarial actions. These coordinated strategies can amplify individual malicious impact through strategic cooperation, creating correlated deviations that may appear legitimate to individual clients with limited neighborhood views.
The fundamental challenge for DFL robustness lies in defending against these varied attack strategies under graph-limited visibility, where each client must make aggregation decisions based solely on local neighborhood information rather than global network state. This constraint makes similarity-based filtering particularly attractive, as it provides a principled approach to neighbor screening that remains effective across different attack types while being computationally tractable for individual clients.
2.3. Byzantine-Robust DFL Defenses
As explained in the previous section, existing DFL defenses have converged on similarity-based neighbor filtering as the dominant paradigm, with methods differing primarily in their filtering criteria and aggregation mechanisms.
UBAR (Guo et al., 2021) utilizes two-stage neighbor selection based on distance and loss filtering, followed by averaging.
LEARN (El-Mhamdi et al., 2021) employs multiple rounds of model exchanges per iteration with trimmed-mean aggregation. The multi-round communication exacerbates the pre-existing scalability issues with similarity-based filtering mechanisms.
SCCLIP (He et al., 2022) applies self-centered clipping to each received neighbor model, mitigating large-magnitude deviations. It provides convergence guarantees in non-convex settings but also requires full model comparisons.
BALANCE (Fang et al., 2024) introduces adaptive similarity thresholds for neighbor screening with strong theoretical convergence guarantees in both strongly convex and non-convex settings. It achieves state-of-the-art robustness but at a high computational cost.
WFAgg (Cajaraville-Aboy et al., 2024) proposes multi-filter approaches for dynamic topologies, combining several screening mechanisms while maintaining the same computational complexity limitations.
However, all these methods require computation of similarities between complete $d$-dimensional model vectors with all neighbors, creating bottlenecks that prevent practical web-scale deployments. Table 1 compares SketchGuard with these existing DFL defense approaches.
Table 1. Comparison of Byzantine-robust decentralized aggregation algorithms. “Convex”/“Non-convex” indicate convergence guarantees; “No $c_{i}$” means no need to know compromised node ratio; “No Complete” means no complete-graph assumption.
Algorithm Convex Non-convex No $c_{i}$ No Complete Scalable
UBAR (Guo et al., 2021) Med.
LEARN (El-Mhamdi et al., 2021) Low
SCCLIP (He et al., 2022) Med.
BALANCE (Fang et al., 2024) Med.
WFAgg (Cajaraville-Aboy et al., 2024) Med.
SketchGuard High
2.4. Compression Techniques in FL
Communication and computational efficiency remain fundamental challenges in FL, motivating extensive research in model compression. Quantization techniques reduce parameter precision (Alistarh et al., 2017; Bernstein et al., 2018), sparsification methods transmit only significant updates (Stich et al., 2018; Sattler et al., 2019), and low-rank approaches exploit model structure (Konečnỳ et al., 2016; Haddadpour et al., 2021). However, most compression techniques target benign settings and often compromise robustness when Byzantine participants are present. Traditional quantization can amplify malicious update impact, while sparsification enables attackers to concentrate influence in transmitted coordinates. More broadly, existing compression methods are incompatible with Byzantine-robust aggregation algorithms, preventing scalable deployment of robust FL.
2.5. Count Sketch
Count Sketch (Charikar et al., 2002) provides randomized linear projection with properties that make it particularly suitable for Byzantine-robust neighbor screening in DFL. Given $\mathbf{w}\in\mathbb{R}^{d}$, a Count Sketch of size $k\ll d$ is constructed using hash function $h:[d]\rightarrow[k]$ and sign function $s:[d]\rightarrow\{-1,+1\}$:
(4) $\mathrm{CS}(\mathbf{w})[b]=\sum_{i:h(i)=b}s(i)w_{i},\quad b=1,\ldots,k.$
Count Sketch possesses three critical properties for robust DFL screening. First, linearity ensures consistent compression across clients, as $\mathrm{CS}(\alpha\mathbf{u}+\beta\mathbf{v})=\alpha\mathrm{CS}(\mathbf{u})+\beta\mathrm{CS}(\mathbf{v})$. Second, unbiased coordinate estimation provides formal recovery guarantees (Charikar et al., 2002; Rothchild et al., 2020). Third, and most crucial for similarity-based Byzantine filtering, Count Sketch preserves approximate Euclidean distances:
Lemma 0 (Distance Preservation (Charikar et al., 2002)).
For any vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{d}$ and sketch size $k=O(\epsilon^{-2}\log(1/\delta))$, with probability at least $1-\delta$:
(5) $(1-\epsilon)\|\mathbf{u}-\mathbf{v}\|^{2}\leq\|\mathrm{CS}(\mathbf{u})-\mathrm{CS}(\mathbf{v})\|^{2}\leq(1+\epsilon)\|\mathbf{u}-\mathbf{v}\|^{2}$
3. SketchGuard: Scalable Robust Aggregation
We now present SketchGuard, which leverages Lemma 1 to perform neighbor screening in the sketch domain. The key design rationale is that if Count Sketch approximately preserves distances, then similarity-based filtering decisions made using sketches will closely match those made using full-precision models, allowing us to defer expensive full-model exchanges until after filtering.
3.1. Protocol Description
Figure 1. The SketchGuard Protocol:
At each DFL training round $t$, client $i$ executes the four-phase SketchGuard protocol detailed in Algorithm 1 and illustrated in Figure 1 as follows:
Phase 1: Local Training. Client $i$ performs local stochastic gradient descent as shown in Equation 2.
Phase 2: Sketch Exchange. The updated model is compressed using CS: $\mathbf{s}_{i}^{t+1/2}=\mathrm{CS}(\mathbf{w}_{i}^{t+1/2})$, and these $k$-dimensional sketches are exchanged with the immediate neighbors $\mathcal{N}_{i}$.
Phase 3: Adaptive Filtering. Neighbor $j$ is accepted if their sketch distance satisfies:
(6) $\|\mathbf{s}_{i}^{t+1/2}-\mathbf{s}_{j}^{t+1/2}\|\leq\gamma\exp(-\kappa t/T)\|\mathbf{s}_{i}^{t+1/2}\|$
where $\gamma$ controls the base threshold, $\kappa$ the decay rate, and $T$ the total rounds. This adaptive threshold mechanism is adopted from BALANCE (Fang et al., 2024), where the exponential decay reflects convergence of honest clients over time.
Remark 1.
When using Count Sketch compression with approximation parameter $\epsilon$, the effective threshold parameter becomes $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$ to account for distance preservation errors in the sketch domain, as established by the analysis in Lemma 1. This ensures that our theoretical convergence guarantees in Section 4 accurately reflect the impact of compression on filtering decisions.
Phase 4: Model Aggregation. Full models are fetched from accepted neighbors $\mathcal{S}_{i}^{t}$. Before aggregation, each received model $\mathbf{w}_{j}^{t+1/2}$ is verified by recomputing its sketch and comparing with the originally received $\mathbf{s}_{j}^{t+1/2}$. Any neighbor whose model fails verification is removed from $\mathcal{S}_{i}^{t}$. The verified models are then aggregated:
(7) $\mathbf{w}_{i}^{t+1}=\alpha\mathbf{w}_{i}^{t+1/2}+\frac{(1-\alpha)}{|\mathcal{S}_{i}^{t}|}\sum_{j\in\mathcal{S}_{i}^{t}}\mathbf{w}_{j}^{t+1/2}$
where $\alpha\in[0,1]$ balances self-reliance and collaboration.
Algorithm 1 SketchGuard: Robust Aggregation via Adaptive Sketch-Based Filtering
0:  Local data $\mathcal{D}_{i}$, neighbors $\mathcal{N}_{i}$, parameters $\gamma,\kappa,\alpha$, sketch size $k$
0:  Updated model $\mathbf{w}_{i}^{t+1}$
1:  $\mathbf{w}_{i}^{t+1/2}\leftarrow\mathbf{w}_{i}^{t}-\eta\mathbf{g}(\mathbf{w}_{i}^{t})$
2:  $\mathbf{s}_{i}^{t+1/2}\leftarrow\mathrm{CS}(\mathbf{w}_{i}^{t+1/2})$
3:  Exchange sketches $\mathbf{s}_{i}^{t+1/2}$ with neighbors $\mathcal{N}_{i}$
4:  $\tau\leftarrow\gamma\exp(-\kappa t/T)\|\mathbf{s}_{i}^{t+1/2}\|$
5:  $\mathcal{S}_{i}^{t}\leftarrow\{j\in\mathcal{N}_{i}:\|\mathbf{s}_{i}^{t+1/2}-\mathbf{s}_{j}^{t+1/2}\|\leq\tau\}$
6:  if $|\mathcal{S}_{i}^{t}|=0$ and $|\mathcal{N}_{i}|>0$ then
7:   $\mathcal{S}_{i}^{t}\leftarrow\{\arg\min_{j\in\mathcal{N}_{i}}\|\mathbf{s}_{i}^{t+1/2}-\mathbf{s}_{j}^{t+1/2}\|\}$
8:  end if
9:  Fetch models $\{\mathbf{w}_{j}^{t+1/2}\}_{j\in\mathcal{S}_{i}^{t}}$ from accepted neighbors
10:  Verify $\mathrm{CS}(\mathbf{w}_{j}^{t+1/2})=\mathbf{s}_{j}^{t+1/2}$ for each $j\in\mathcal{S}_{i}^{t}$; remove if mismatch
11:  $\mathbf{w}_{i}^{t+1}\leftarrow\alpha\mathbf{w}_{i}^{t+1/2}+\frac{1-\alpha}{|\mathcal{S}_{i}^{t}|}\sum_{j\in\mathcal{S}_{i}^{t}}\mathbf{w}_{j}^{t+1/2}$
12:  return $\mathbf{w}_{i}^{t+1}$
3.2. Complexity Analysis and Performance Trade-offs
We analyze the per-node-per-round complexity of SketchGuard compared to the best-case computation and communication complexity of existing Byzantine-robust methods. Let $d$ denote the model dimension, $|\mathcal{N}_{i}|$ the number of neighbors, $k$ the sketch size, and $|\mathcal{S}_{i}^{t}|$ the number of accepted neighbors. For computational complexity, SketchGuard executes four phases per round:
  • Sketch generation: $O(d)$
  • Neighbor screening (compressed domain): $O(k\cdot|\mathcal{N}_{i}|)$
  • Model verification & aggregation (accepted neighbors): $O(d\cdot|\mathcal{S}_{i}^{t}|)$
This yields total complexity $O(d+k\cdot|\mathcal{N}_{i}|+d\cdot|\mathcal{S}_{i}^{t}|)$, compared to $O(d\cdot|\mathcal{N}_{i}|)$ for traditional methods that perform full-precision screening of all neighbors.
Table 2. Computational complexity comparison with state-of-the-art (SOTA) across training phases.
Training Phase SOTA SketchGuard
Local Training $O(d)$ $O(d)$
Sketch Generation $O(d)$
Neighbor Screening $O(d\cdot|\mathcal{N}_{i}|)$ $O(k\cdot|\mathcal{N}_{i}|)$
Model Verification & Aggregation $O(d\cdot|\mathcal{N}_{i}|)$ $O(d\cdot|\mathcal{S}_{i}^{t}|)$
Total Per Round $O(d\cdot|\mathcal{N}_{i}|)$ $O(d+k\cdot|\mathcal{N}_{i}|+d\cdot|\mathcal{S}_{i}^{t}|)$
For communication complexity, SketchGuard transmits two types of messages: sketches to all neighbors and full models to accepted neighbors. This results in $O(k\cdot|\mathcal{N}_{i}|+d\cdot|\mathcal{S}_{i}^{t}|)$ parameters per round, compared to $O(d\cdot|\mathcal{N}_{i}|)$ for existing approaches.
The theoretical requirement $k=O(\epsilon^{-2}\log(1/\delta))$ translates to practical sketch sizes that remain much smaller than model dimensions while providing strong approximation guarantees. SketchGuard’s performance benefits scale with three factors: (1) the compression ratio $d/k$, where larger models yield greater savings, (2) the filtering effectiveness ratio $|\mathcal{S}_{i}^{t}|/|\mathcal{N}_{i}|$, where successful Byzantine rejection reduces the number of full models fetched, and (3) network degree $|\mathcal{N}_{i}|$, where denser connectivity amplifies the screening overhead reduced by compression. While sketch generation and exchange introduce $O(d+k\cdot|\mathcal{N}_{i}|)$ overhead even in benign settings, this cost becomes negligible as model dimensionality and network scale increase, which characterizes the web-scale deployments where Byzantine robustness is necessary.
4. Convergence Analysis
In this section, we establish that SketchGuard maintains convergence guarantees despite compressed filtering by proving that sketch-based screening preserves the robustness properties of full-precision similarity-based defenses with controlled degradation bounds.
4.1. Technical Assumptions
We state the standard assumptions that impact on convergence analysis.
Assumption 1 (Strong Convexity).
The population risk $F(\mathbf{w})$ is $\mu$-strongly convex, i.e., for all $\mathbf{w}_{1},\mathbf{w}_{2}\in\Theta$, one has that:
$F(\mathbf{w}_{1})+\langle\nabla F(\mathbf{w}_{1}),\mathbf{w}_{2}-\mathbf{w}_{1}\rangle+\frac{\mu}{2}\|\mathbf{w}_{2}-\mathbf{w}_{1}\|^{2}\leq F(\mathbf{w}_{2}).$
Assumption 2 (Smoothness).
The population risk $F(\mathbf{w})$ is $L$-smooth, i.e., for all $\mathbf{w}_{1},\mathbf{w}_{2}\in\Theta$, we have that:
$\|\nabla F(\mathbf{w}_{1})-\nabla F(\mathbf{w}_{2})\|\leq L\|\mathbf{w}_{1}-\mathbf{w}_{2}\|.$
Assumption 3 (Bounded Variance).
The stochastic gradient $\mathbf{g}(\mathbf{w}_{i})$ computed by an honest client $i\in\mathcal{H}$ is an unbiased estimator of the true gradient, and $\mathbf{g}(\mathbf{w}_{i})$ has bounded variance, where $\mathcal{H}$ is the set of honest clients. That is, $\forall i\in\mathcal{H}$, one has that:
$\mathbb{E}[\mathbf{g}(\mathbf{w}_{i})]=\nabla F(\mathbf{w}_{i}),\quad\mathbb{E}[\|\mathbf{g}(\mathbf{w}_{i})-\nabla F(\mathbf{w}_{i})\|^{2}]\leq\delta^{2}.$
Assumption 4 (Bounded Parameters).
For any honest client $i\in\mathcal{H}$, the model $\mathbf{w}_{i}$ and $\|\nabla F(\mathbf{w}_{i})\|$ are bounded. That is, $\forall i\in\mathcal{H}$, we have $\|\mathbf{w}_{i}\|\leq\psi$, and $\|\nabla F(\mathbf{w}_{i})\|\leq\rho$.
Assumption 5 (Graph Connectivity).
The subgraph induced by honest clients $G_{\mathcal{H}}$ remains connected throughout training.
Assumption 6 (Hash Function Synchronization).
All clients use identical hash and sign functions for consistent sketching across the network.
4.2. Main Convergence Results
Leveraging Lemma 1, we establish convergence guarantees for SketchGuard that achieve similar results to those based on full-precision methods.
Theorem 1 (Strongly Convex Convergence with Compression).
For $\mu$-strongly convex and $L$-smooth objectives, with learning rate $\eta\leq\min\{1/(4L),1/\mu\}$ and effective threshold parameter $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$, after $T$ rounds:
$\mathbb{E}[F(\mathbf{w}_{i}^{T})-F(\mathbf{w}^{*})]\leq(1-\mu\eta)^{T}[F(\mathbf{w}_{i}^{0})-F(\mathbf{w}^{*})]$
(8) $\quad+\frac{2L\eta\delta^{2}}{\mu}+\frac{2\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\mu\eta}$
where the compression error manifests only through the $(1+\epsilon)/(1-\epsilon)$ factor in $\gamma_{\text{eff}}$.
Proof.
See Appendix A. ∎
Theorem 2 (Non-Convex Convergence with Compression).
For non-convex $L$-smooth objectives, with the same parameter choices:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}]\leq\frac{2[F(\mathbf{w}_{i}^{0})-F^{*}]}{\eta T}$
(9) $\quad+4L\eta\delta^{2}+\frac{4\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\eta}$
Proof.
See Appendix B. ∎
Both theorems demonstrate that SketchGuard achieves optimal convergence rates for their respective settings (Garrigos and Gower, 2023), with sketch approximation introducing only a $(1+O(\epsilon))$ multiplicative factor on the threshold-dependent terms in the convergence bounds. For practical values of $\epsilon$ (e.g., $\epsilon=0.1$ yields $\gamma_{\text{eff}}\approx 1.1\gamma$), this degradation is minimal while enabling substantial compression.
4.3. Robustness Preservation Under Compression
Our final theoretical result establishes that sketch compression preserves Byzantine robustness guarantees with controlled degradation.
Lemma 0 (Sketch Compression Preserves Byzantine Resilience).
Under the conditions of Theorems 1 and 2, sketch-based filtering with effective threshold $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$ provides equivalent Byzantine resilience to full-precision filtering with threshold $\gamma_{\text{eff}}$. Specifically, compression does not create new attack vectors beyond the controlled threshold degradation.
Proof.
This follows directly from the convergence analysis in Theorems 1 and 2. Both proofs bound the neighbor difference terms under sketch-based filtering (Appendix A.5):
$\left\|\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|\leq\gamma_{\text{eff}}\psi$
This bound is identical in form to the full-precision case, differing only in the threshold parameter $\gamma\to\gamma_{\text{eff}}$. Since the convergence proofs do not require any additional structural assumptions about the filtering mechanism beyond this bound, any Byzantine strategy must satisfy the same constraints as in the full-precision case, but with the relaxed threshold $\gamma_{\text{eff}}$. Therefore, compression preserves the Byzantine resilience properties established by (Fang et al., 2024), with degradation bounded by the factor $\sqrt{(1+\epsilon)/(1-\epsilon)}=1+O(\epsilon)$. ∎
5. Performance Evaluation
We evaluate SketchGuard [^1] Code and experiment artifacts available at doi.org/10.5281/zenodo.17223405. through comprehensive experiments demonstrating that sketch-based compression maintains identical Byzantine robustness to state-of-the-art full-precision defenses while substantially reducing computational and communication overhead.
5.1. Experimental Setup
5.1.1. Datasets and Models
We conduct experiments on two standard federated learning benchmarks from the LEAF suite (Caldas et al., 2018), chosen for their naturally non-IID data distributions that closely reflect real-world heterogeneity in federated environments. We evaluate both robustness and scalability across diverse model sizes.
FEMNIST (Federated Extended MNIST) is a character recognition dataset with 62 classes distributed across 3,550 users, where each user represents a different writer’s handwriting samples. We employ a convolutional neural network comprising two convolutional layers followed by two fully connected layers. Our baseline configuration uses 6.6M parameters, which we systematically scale from 220K to over 60M parameters in our dimensionality scaling experiments (Section 5.3) to evaluate SketchGuard under varying model complexities representative of modern federated learning deployments.
CelebA is a celebrity face attributes dataset distributed across 9,343 users. We use it for binary smile classification (smiling or not smiling), with a CNN architecture adapted for RGB inputs ($84\times 84$ pixels) containing 2.2M parameters. The network consists of two convolutional layers and two fully connected layers. Complete architectural specifications for both models are provided in Appendix C.1.
5.1.2. Network Topologies
(a) Ring(b) ER Sparsep=0.2p=0.2(c) ER Mediump=0.45p=0.45(d) ER Densep=0.6p=0.6(e) Fully Connected
Figure 2. Network topologies used in the robustness evaluation experiments.
We evaluate defense mechanisms across carefully selected network topologies that capture the spectrum of connectivity patterns in real-world decentralized systems, from resource-constrained sparse networks to well-connected infrastructures. For robustness evaluation experiments in Section 5.2, we employ Erdős-Rényi (ER) random graphs (Erdős and Rényi, 1960), which naturally model peer-to-peer networks where connections form probabilistically. ER graphs are generated by independently connecting each pair of nodes with probability $p$, producing networks with variable local connectivity that reflect real-world decentralized learning scenarios where participants discover neighbors organically rather than through engineered topologies. We evaluate five topologies spanning from minimal connectivity (Ring) to maximum connectivity (Fully Connected), with three intermediate ER graph densities as shown in Figure 2; complete specifications are provided in Appendix C.2.1. These experiments use 20 nodes to enable comprehensive evaluation across varying malicious client fractions and defense mechanisms while maintaining computational tractability. For network scaling experiments in Section 5.3, we employ k-regular graphs (Bondy et al., 1976) where each node has exactly $k$ neighbors. This design choice ensures consistent per-node computational load across all defense mechanisms, enabling fair comparison of algorithmic efficiency independent of topology-induced variance. We evaluate networks scaling from 20 to 300 nodes, demonstrating SketchGuard’s scalability to large-scale deployments with hundreds of participants.
5.1.3. Attack Models
We evaluate Byzantine robustness using two representative attack strategies that represent different threat models and attack sophistication levels.
Directed Deviation attacks employ an optimization-based approach that crafts malicious updates by solving for parameters that maximize deviation in the direction opposite to honest gradient descent. This adaptive attack strategy represents a strong adversary that actively attempts to subvert Byzantine-robust aggregation mechanisms by carefully positioning malicious updates to evade detection while maximizing impact. We implement the directed deviation attack following the formulation in (Fang et al., 2020), which has demonstrated effectiveness against common Byzantine-robust aggregation rules including Krum and trimmed mean.
Gaussian attacks (Blanchard et al., 2017) inject random noise $\hat{\mathbf{w}}_{j}=\mathbf{w}_{j}+\mathcal{N}(0,\sigma^{2}\mathbf{I})$ with $\sigma=1$, representing less sophisticated but realistic adversaries that disrupt convergence through stochastic perturbations. While simpler than directed deviation, these attacks model practical scenarios where attackers lack complete knowledge of the aggregation mechanism or resources for optimization-based attacks.
We vary the fraction of Byzantine clients from 0% to 80% in 10% increments to evaluate robustness across different attack intensities. These attack strategies provide complementary evaluation: directed deviation establishes defense performance against sophisticated adaptive adversaries, while Gaussian attacks validate robustness against simpler but more common threat models.
5.1.4. Baselines and Configuration
We compare SketchGuard against four baseline methods: D-FedAvg (Lian et al., 2017), KRUM (Blanchard et al., 2017), UBAR (Guo et al., 2021), and BALANCE (Fang et al., 2024), representing the spectrum from non-robust to state-of-the-art Byzantine-robust aggregation. For all experiments, except the $k$-ablation study, we employ sketch sizes $k=1000$ for FEMNIST and $k=350$ for CelebA. For the baseline model configurations, this achieves compression ratios of 6603:1 and 6342:1 respectively, with approximation error $\epsilon\lesssim 0.2$ and probability $1-\delta>0.99$. All experiments use standard DFL configurations with 10 global rounds and 3 local epochs per round. We repeat each experiment across 3 random seeds and report mean results; observed variances are minimal across all metrics. Complete hyperparameter specifications are provided in Appendix C.2.2.
5.1.5. Evaluation Metrics
We evaluate defense mechanisms using three complementary metrics that capture robustness, computational efficiency, and communication overhead.
Test Error Rate (TER). Defined as 1test accuracy1-$\text${test accuracy}, TER measures performance degradation averaged across honest clients. Lower values indicate better Byzantine resilience.
Per-Round Computation Time. Wall-clock time for neighbor screening and aggregation phases per client per round, excluding local training which remains constant across methods. All measurements are recorded on identical hardware.
Communication Overhead. Total parameters transmitted per client per round, including sketches and full model exchanges. We report both absolute counts and reduction percentages relative to full-precision baselines.
5.2. Byzantine Robustness Evaluation
5.2.1. Attack Strategy Resilience
Figure 3. Impact of fraction of malicious clients on TER across different datasets and attack types.
Figure 3 evaluates SketchGuard’s robustness against both directed deviation and Gaussian attacks on FEMNIST and CelebA datasets, averaged across all network topologies. SketchGuard achieves robustness equivalent to state-of-the-art methods BALANCE and UBAR across both attack strategies. Compared to BALANCE, SketchGuard exhibits statistically identical performance with mean absolute TER deviation of 0.02 percentage points and maximum deviation of 0.49 percentage points. Compared to UBAR, SketchGuard shows mean absolute TER deviation of 0.35 percentage points and maximum deviation of 1.51 percentage points, with SketchGuard having marginally higher TER on average by 0.20 percentage points; however, UBAR’s loss-based screening mechanism requires significantly higher computational overhead than both BALANCE and SketchGuard, as demonstrated in Section 5.3. Against both directed deviation (Figures 3a, 3c) and Gaussian attacks (Figures 3b, 3d), SketchGuard maintains consistently low error rates across all Byzantine client fractions, demonstrating that sketch-based compression preserves the robustness properties of full-precision similarity-based defenses while enabling substantial efficiency gains.
5.2.2. Sketch Size Robustness
SketchGuard demonstrates remarkable robustness to sketch size selection. Across $k$ values ranging from 500 to 100,000 on both FEMNIST and CelebA datasets under 50% Byzantine clients with directed deviation attacks, performance remains completely stable (TER = 15.59% for FEMNIST, 10.35% for CelebA across all $k$ values). Even with compression ratios exceeding 13,000:1 ($k=500$ for FEMNIST’s 6.6M parameters), SketchGuard preserves full Byzantine resilience. This insensitivity to $k$ stems from Count Sketch’s distance preservation guarantees: while smaller $k$ values introduce larger approximation errors, the resulting sketch-based distances remain sufficiently accurate to distinguish Byzantine node signatures from honest node behavior. The adaptive threshold mechanism $\gamma_{\text{eff}}$ compensates for increased approximation error, ensuring robust filtering across the entire practical range of sketch sizes.
5.2.3. Topology-Dependent Robustness
Figure 4. Impact of fraction of malicious clients on TER across different network topologies.
Figure 4 presents SketchGuard’s robustness across five network topologies, averaged over attack types and datasets. SketchGuard demonstrates identical robustness to BALANCE and UBAR across all network topologies, from minimal connectivity in Ring topology to maximum connectivity in Fully Connected networks. In well-connected topologies (ER graphs with $p\geq 0.45$, Fully Connected), all three Byzantine-robust methods maintain low error rates below 20% even when 80% of the network is compromised, significantly outperforming KRUM and the non-robust D-FedAvg baseline. In extremely sparse topologies (Ring), BALANCE, UBAR, and SketchGuard exhibit graceful degradation, tolerating up to 40% Byzantine clients before substantial performance loss. Despite having only two neighbors per node in the Ring topology, SketchGuard maintains performance parity with full-precision methods across all attack intensities, demonstrating that sketch-based compression does not compromise filtering effectiveness even under severe connectivity constraints.
5.3. Computational Efficiency
Figure 5 evaluates SketchGuard’s computational efficiency through network size and model dimensionality scaling experiments on FEMNIST with 50% Byzantine clients under directed deviation attacks. Network scaling uses the baseline 6.6M parameter model across varying k-regular graph sizes, while model dimensionality scaling uses a fixed k-regular graph with node degree 154 across varying model sizes as detailed in Appendix C.3.
Figure 5. Impact of network size and model dimensionality on per-node computation time.
5.3.1. Network Size Scalability
SketchGuard exhibits near-constant computational cost as network connectivity increases from 16 to 299 neighbors (Figure 5a), averaging 0.35s per round, while BALANCE and UBAR scale linearly with neighbor count. At smaller network sizes ($\leq$32 neighbors), SketchGuard’s sketch generation overhead dominates, resulting in higher absolute costs than full-precision methods. However, as connectivity increases beyond 96 neighbors, SketchGuard’s compressed screening operations yield substantial benefits. The efficiency gap widens multiplicatively with network size: at 154 neighbors, SketchGuard requires 0.34s compared to BALANCE’s 0.51s and UBAR’s 1.10s, achieving 33% and 69% savings respectively; at 299 neighbors, these savings expand to 60% and 82%, with SketchGuard requiring only 0.39s versus BALANCE’s 0.99s and UBAR’s 2.14s. This scalability advantage stems from neighbor screening operating in compressed $k$-dimensional space rather than full $d$-dimensional space, yielding multiplicative gains proportional to network connectivity.
5.3.2. Model Dimensionality Scalability
SketchGuard demonstrates sub-linear scaling as model size increases from 220K to 60M parameters (Figure 5b), contrasting sharply with the linear scaling of BALANCE and UBAR. For models exceeding 26M parameters, SketchGuard maintains average computational improvement of 64% over state-of-the-art methods. At 60M parameters, SketchGuard requires 2.0s versus BALANCE’s 4.2s and UBAR’s 10.0s, achieving 52% and 80% reductions respectively. This sub-linear scaling emerges from Count Sketch’s property that sketch size $k$ depends on approximation parameters ($\epsilon$, $\delta$) rather than model dimension $d$. While full-precision methods incur $O(d|\mathcal{N}_{i}|)$ comparison costs that grow linearly with model size, SketchGuard’s sketch comparison cost remains fixed at $O(k|\mathcal{N}_{i}|)$ regardless of $d$, proving particularly valuable for large-scale DFL scenarios involving billion-parameter models (Borzunov et al., 2023).
5.4. Communication Efficiency
SketchGuard reduces per-round communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$, where $|\mathcal{S}_{i}|\leq|\mathcal{N}_{i}|$ denotes accepted neighbors post-filtering.
5.4.1. Compression Ratios and Bandwidth Reduction
With compression ratios exceeding 6300:1 for both datasets, sketch transmission overhead becomes negligible compared to full model exchanges. Table 3 presents communication efficiency under different filtering scenarios. In scenarios where Byzantine filtering successfully rejects 70% of malicious neighbors, SketchGuard achieves approximately 70% communication reduction relative to full-precision methods. Even in benign settings where all neighbors are accepted, the sketch overhead of $O(k|\mathcal{N}_{i}|)$ represents less than 0.02% additional cost compared to the $O(d|\mathcal{N}_{i}|)$ baseline due to the compression ratio $k/d\approx 1/6300$, introducing negligible bandwidth penalty.
Table 3. Communication Overhead Comparison
Scenario Accepted Neighbors Reduction
Benign (no filtering) $|\mathcal{S}_{i}|=|\mathcal{N}_{i}|$ <<0.02% penalty
50% Byzantine filtering $|\mathcal{S}_{i}|\approx 0.5|\mathcal{N}_{i}|$ $\sim$50%
70% Byzantine filtering $|\mathcal{S}_{i}|\approx 0.3|\mathcal{N}_{i}|$ $\sim$70%
5.4.2. Multiplicative Scaling Benefits
The communication savings scale multiplicatively with both model dimensionality and network connectivity. Consider a fully connected network with 100 neighbors per node and FEMNIST’s 6.6M parameter model under 50% Byzantine clients, where Count Sketch compresses models to 1K parameters:
  • SketchGuard: Transmits 100K sketch parameters plus 330M full model parameters, totaling $\sim$330M parameters.
  • Full-Precision Methods: Transmit 660M parameters to all neighbors.
  • Reduction: $\sim$50% bandwidth savings.
These bandwidth savings prove critical for resource-constrained devices and bandwidth-limited networks characteristic of real-world decentralized learning deployments. As model sizes grow and networks expand, SketchGuard’s communication efficiency becomes essential for practical deployment.
6. Conclusions and Future Work
We proposed SketchGuard, a framework for scaling Byzantine-robust decentralized federated learning through sketch-based neighbor screening. By decoupling Byzantine filtering from model aggregation, our approach reduces communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$ while preserving robustness guarantees of full-precision similarity-based defenses.
Our theoretical analysis establishes that Count Sketch compression maintains Byzantine resilience with controlled degradation bounds, where approximation errors introduce only a $(1+O(\epsilon))$ factor in the effective threshold parameter. We prove convergence guarantees in both strongly convex and non-convex settings, achieving optimal rates for their respective problem classes. Comprehensive experiments demonstrate that SketchGuard maintains identical robustness to state-of-the-art methods, with mean TER deviation of only up to 0.35 percentage points. This robustness equivalence is achieved while reducing computation time by up to 82%, with efficiency gains scaling multiplicatively with model dimensionality and network connectivity. Communication overhead scales proportionally with the fraction of accepted neighbors, achieving reductions of 50-70% under typical Byzantine attack scenarios.
SketchGuard addresses the fundamental scalability bottleneck limiting real-world deployment of robust decentralized learning. Future work includes efficient hash synchronization protocols for dynamic networks, adaptive sketch sizing based on attack intensity, and theoretical extensions to non-IID data distributions. As FL expands into web-scale deployments with untrusted participants, scalable robustness mechanisms like SketchGuard become essential infrastructure for maintaining both security and performance in distributed machine learning systems.
Acknowledgements.
This work is supported by the Australian Research Council (ARC) through Discovery Project grant DP240102088 and Linkage Infrastructure, Equipment and Facilities grant LE200100049.
Appendix A Proof of Theorem 1
We provide a detailed proof of the convergence guarantee for SketchGuard under strongly convex objectives with sketch-based compression. The proof structure follows the analysis framework of BALANCE (Fang et al., 2024), with key modifications to handle Count Sketch compression errors through the effective threshold parameter $\gamma_{\text{eff}}$.
A.1. Setup and Notation
We denote by $\mathbf{g}(\mathbf{w}_{i}^{t})$ the stochastic gradient computed by client $i$ at round $t$. The local update is:
(10) $\mathbf{w}_{i}^{t+1/2}=\mathbf{w}_{i}^{t}-\eta\mathbf{g}(\mathbf{w}_{i}^{t}),$
where $\eta>0$ is the learning rate. For simplicity in the proof, we drop the superscript $t$ in $\mathcal{S}_{i}^{t}$ and write $\mathcal{S}_{i}$.
A.2. Model Update Analysis
The model update for client $i$ can be expressed as:
$\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}$ $=\alpha\mathbf{w}_{i}^{t+1/2}$
(11) $\quad+(1-\alpha)\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t}$
$=\alpha\mathbf{w}_{i}^{t+1/2}+(1-\alpha)\frac{1}{|\mathcal{S}_{i}|}$
(12) $\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2}+\mathbf{w}_{i}^{t+1/2})-\mathbf{w}_{i}^{t}$
$=\mathbf{w}_{i}^{t+1/2}+\frac{1-\alpha}{|\mathcal{S}_{i}|}$
(13) $\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})-\mathbf{w}_{i}^{t}$
$=-\eta\mathbf{g}(\mathbf{w}_{i}^{t})+\frac{1-\alpha}{|\mathcal{S}_{i}|}$
(14) $\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})$
A.3. Applying Smoothness
By the $L$-smoothness assumption, we have:
$F(\mathbf{w}_{i}^{t+1})$ $\leq F(\mathbf{w}_{i}^{t})+\langle\nabla F(\mathbf{w}_{i}^{t}),\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\rangle$
(15) $\quad+\frac{L}{2}\|\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\|^{2}$
Substituting the model update expression:
$F(\mathbf{w}_{i}^{t+1})$ $\leq F(\mathbf{w}_{i}^{t})-\eta\langle\nabla F(\mathbf{w}_{i}^{t}),\mathbf{g}(\mathbf{w}_{i}^{t})\rangle$
$\quad+\langle\nabla F(\mathbf{w}_{i}^{t}),\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\rangle$
$\quad+\frac{L}{2}\left\|-\eta\mathbf{g}(\mathbf{w}_{i}^{t})\right$.
(16) $\quad\left.+\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$
A.4. Bounding the Quadratic Term
Using the inequality $\|a+b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$:
$\left\|-\eta\mathbf{g}(\mathbf{w}_{i}^{t})+\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$
$\leq 2\eta^{2}\|\mathbf{g}(\mathbf{w}_{i}^{t})\|^{2}$
(17) $\quad+2\left\|\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$
A.5. Impact of Sketch-Based Filtering
With sketch-based filtering, neighbor $j$ is accepted if:
$\|\mathrm{CS}(\mathbf{w}_{i}^{t+1/2})-\mathrm{CS}(\mathbf{w}_{j}^{t+1/2})\|$
(18) $\quad\leq\gamma\exp(-\kappa t/T)\|\mathrm{CS}(\mathbf{w}_{i}^{t+1/2})\|$
By the distance preservation property of Count Sketch (Charikar et al., 2002), this implies:
$\|\mathbf{w}_{i}^{t+1/2}-\mathbf{w}_{j}^{t+1/2}\|$
(19) $\quad\leq\gamma\sqrt{\frac{1+\epsilon}{1-\epsilon}}\exp(-\kappa t/T)\|\mathbf{w}_{i}^{t+1/2}\|$
Define $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$. Then:
(20) $\left\|\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|\leq\gamma_{\text{eff}}\psi$
where we used the bounded parameters assumption that $\|\mathbf{w}_{i}^{t+1/2}\|\leq\psi$.
A.6. Taking Expectation
Taking expectation and using the bounded variance assumption:
$\mathbb{E}[F(\mathbf{w}_{i}^{t+1})]$ $\leq\mathbb{E}[F(\mathbf{w}_{i}^{t})]-\eta\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}$
$\quad+(1-\alpha)\gamma_{\text{eff}}\psi\rho$
$\quad+L\eta^{2}(\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}+\delta^{2})$
(21) $\quad+L(1-\alpha)^{2}\gamma_{\text{eff}}^{2}\psi^{2}$
A.7. Simplifying with Parameter Constraints
With $\eta\leq 1/(4L)$, we have $L\eta^{2}\leq\eta/4$:
$\mathbb{E}[F(\mathbf{w}_{i}^{t+1})]$ $\leq\mathbb{E}[F(\mathbf{w}_{i}^{t})]-\frac{\eta}{2}\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}$
(22) $\quad+L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
where we chose $\gamma_{\text{eff}}\leq\rho/(L\psi(1-\alpha))$.
A.8. Applying Strong Convexity
By the strong convexity assumption and the Polyak-Łojasiewicz inequality:
(23) $\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}\geq 2\mu(F(\mathbf{w}_{i}^{t})-F(\mathbf{w}^{*}))$
Therefore:
$\mathbb{E}[F(\mathbf{w}_{i}^{t+1})-F(\mathbf{w}^{*})]$
$\quad\leq(1-\mu\eta)\mathbb{E}[F(\mathbf{w}_{i}^{t})-F(\mathbf{w}^{*})]$
(24) $\quad+L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
A.9. Telescoping and Final Bound
Telescoping over $t=0,1,\ldots,T-1$:
$\mathbb{E}[F(\mathbf{w}_{i}^{T})-F(\mathbf{w}^{*})]$
$\quad\leq(1-\mu\eta)^{T}[F(\mathbf{w}_{i}^{0})-F(\mathbf{w}^{*})]$
$\quad+\sum_{t=0}^{T-1}(1-\mu\eta)^{T-1-t}$
$\quad\times(L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha))$
$\quad=(1-\mu\eta)^{T}[F(\mathbf{w}_{i}^{0})-F(\mathbf{w}^{*})]$
$\quad+\frac{1-(1-\mu\eta)^{T}}{\mu\eta}$
$\quad\times(L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha))$
$\quad\leq(1-\mu\eta)^{T}[F(\mathbf{w}_{i}^{0})-F(\mathbf{w}^{*})]$
(25) $\quad+\frac{2L\eta\delta^{2}}{\mu}+\frac{2\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\mu\eta}$
This completes the proof of Theorem 1. The key observation is that the compression error only affects the convergence through $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$, introducing a controllable degradation factor while maintaining the same convergence rate as the state-of-the-art.
Appendix B Proof of Theorem 2
We establish the convergence guarantee for SketchGuard in non-convex settings, adapting the BALANCE analysis framework (Fang et al., 2024) for sketch compression.
B.1. Starting from Smoothness
Following the analysis from Appendix A up to the expectation bound, we have:
$\mathbb{E}[F(\mathbf{w}_{i}^{t+1})]$ $\leq\mathbb{E}[F(\mathbf{w}_{i}^{t})]-\frac{\eta}{2}\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}$
(26) $\quad+2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
B.2. Rearranging for Gradient Norm
Rearranging the inequality:
$\frac{\eta}{2}\mathbb{E}[\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}]$
$\quad\leq\mathbb{E}[F(\mathbf{w}_{i}^{t})-F(\mathbf{w}_{i}^{t+1})]$
(27) $\quad+2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
B.3. Telescoping
Summing from $t=0$ to T1T-1:
$\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}]$
$\quad\leq F(\mathbf{w}_{i}^{0})-\mathbb{E}[F(\mathbf{w}_{i}^{T})]$
(28) $\quad+T(2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha))$
B.4. Averaging and Using Lower Bound
Since $F(\mathbf{w}_{i}^{T})\geq F^{*}$:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}]$
$\quad\leq\frac{2(F(\mathbf{w}_{i}^{0})-F^{*})}{\eta T}$
(29) $\quad+4L\eta\delta^{2}+\frac{4\gamma_{\text{eff}}\psi\rho(1-\alpha)}{\eta}$
This establishes Theorem 2. The convergence rate matches the optimal rate for non-convex optimization, with the compression error appearing only through $\gamma_{\text{eff}}$.
Appendix C Supplementary Experimental Details
This appendix provides comprehensive details about the experimental setup, including model architectures, hyperparameter configurations, network topologies, and scalability experiment specifications used throughout our study.
C.1. Detailed Model Architectures
We provide complete architectural specifications for the two datasets used in our experiments: FEMNIST and CelebA. Both architectures follow standard designs adapted to their respective input dimensions and task requirements.
C.1.1. FEMNIST Architecture
The FEMNIST model follows the original LEAF specification with a convolutional neural network design. The architecture consists of two convolutional layers followed by max pooling and two fully connected layers. Table 4 provides the complete layer-by-layer specification, including parameter counts and output dimensions.
Table 4. FEMNIST Model Architecture Details
Layer Type Parameters Output Size
Input $1\times 28\times 28$
Conv1 Conv2d $32\times(5\times 5\times 1)+32$ $32\times 28\times 28$
Pool1 MaxPool2d $32\times 14\times 14$
Conv2 Conv2d $64\times(5\times 5\times 32)+64$ $64\times 14\times 14$
Pool2 MaxPool2d $64\times 7\times 7$
Flatten $3136$
FC1 Linear $3136\times 2048+2048$ $2048$
FC2 Linear $2048\times 62+62$ $62$
Total 6,603,710
C.1.2. CelebA Architecture
The CelebA model uses a LeNet-style architecture adapted for larger RGB input images ($84\times 84$ pixels). The network has a similar structure to FEMNIST but with modified convolutional layers to handle color images and a binary classification output. The complete architecture is detailed in Table 5.
Table 5. CelebA Model Architecture Details
Layer Type Parameters Output Size
Input $3\times 84\times 84$
Conv1 Conv2d $30\times(3\times 3\times 3)+30$ $30\times 84\times 84$
Pool1 MaxPool2d $30\times 42\times 42$
Conv2 Conv2d $50\times(3\times 3\times 30)+50$ $50\times 42\times 42$
Pool2 MaxPool2d $50\times 21\times 21$
Flatten $22050$
FC1 Linear $22050\times 100+100$ $100$
FC2 Linear $100\times 2+2$ 22
Total 2,219,692
C.2. Complete Experimental Configuration
This section details all hyperparameters and configuration settings used across our experiments to ensure reproducibility.
C.2.1. Network Topology Specifications
We evaluate defense mechanisms across five different network topologies representing varying levels of connectivity. The ring topology represents minimal connectivity (degree 2), while Erdős-Rényi graphs with varying connection probabilities $p$ provide intermediate connectivity levels. The fully connected topology represents maximum connectivity where each node connects to all others. Table 6 specifies the parameters and expected node degrees for each topology in our 20-node experimental networks.
Table 6. Network Topology Configurations
Topology Parameters Expected Degree
Ring 2
Erdős-Rényi (sparse) $p=0.2$ 3.8
Erdős-Rényi (medium) $p=0.45$ 8.55
Erdős-Rényi (dense) $p=0.6$ 11.4
Fully Connected 19
C.2.2. Hyperparameter Settings
Table 7 presents the complete set of hyperparameters used in our experiments, organized by category. Training configuration parameters were selected based on standard federated learning practices. Defense mechanism parameters ($\gamma$, $\kappa$, $\alpha$) were tuned to balance robustness and model utility. For SketchGuard, we set the sketch size $k=1000$ for FEMNIST and $k=350$ for CelebA to get roughly the same factor of compression for both datasets.
Table 7. Complete Hyperparameter Configuration
Parameter Value
Training Configuration
Number of clients 20
Global epochs 10
Local epochs per round 3
Batch size 64
Learning rate 0.01 (FEMNIST), 0.001 (CelebA)
Maximum samples per client 10,000 (FEMNIST), 4,500 (CelebA)
Random seed 987654321, 39573295, 32599368
SketchGuard Parameters
Sketch size ($k$) 1,000 (FEMNIST), 350 (CelebA)
Threshold parameter ($\gamma$) 2.0
Decay parameter ($\kappa$) 1.0
Mixing parameter ($\alpha$) 0.5
Hash seed 42
BALANCE Parameters
Threshold parameter ($\gamma$) 2.0
Decay parameter ($\kappa$) 1.0
Mixing parameter ($\alpha$) 0.5
UBAR Parameters
Robustness parameter ($\rho$) 1.0attack percentage1.0-$\text${attack percentage}
KRUM Parameters
Compromised fraction Attack percentage
C.3. Scalability Experiments
We conducted two types of scalability experiments to evaluate how defense mechanisms perform under different scaling conditions: network size scaling and model dimensionality scaling.
C.3.1. Network Size Scaling
To evaluate the computational scalability of defense mechanisms with respect to network size, we conducted experiments on k-regular graphs with varying numbers of participants. Table 8 shows the configurations used, with network sizes ranging from 20 to 300 nodes. All networks maintain 50% Byzantine nodes to test defense robustness under high attack scenarios. The node degree was selected to ensure connectivity while maintaining realistic peer-to-peer network constraints. These experiments used shorter training runs (3 rounds with 1 local epoch each) to focus on measuring computational overhead rather than convergence behavior.
Table 8. Network Size Scaling Configurations
Node Degree Network Size Attack Percentage
16 20 50%
32 35 50%
96 100 50%
154 155 50%
299 300 50%
C.3.2. Model Scaling Variants
To investigate the impact of model dimensionality on defense mechanism performance, we created five variants of the FEMNIST architecture with different parameter counts. Table 9 describes these variants, ranging from the Tiny model with approximately 220K parameters to the XLarge model with over 60M parameters. These variants were created by systematically scaling the number of convolutional filters and fully connected layer dimensions while maintaining the overall architectural structure.
Table 9. Model Scaling Variants
Variant Architecture Modifications Parameters
Tiny Reduced filter counts and hidden units 220,318
Small Standard configuration 848,382
Standard Baseline FEMNIST architecture 6,603,710
Large Increased filter counts and hidden units 26,154,814
XLarge Further increased dimensions 60,271,678

### v3

@@
 SketchGuard: Scaling Byzantine-Robust Decentralized Federated Learning via Sketch-Based Screening
-Abstract.
-Decentralized Federated Learning (DFL) enables privacy-preserving collaborative training without centralized servers, but remains vulnerable to Byzantine attacks where malicious clients submit corrupted model updates.
-Existing Byzantine-robust DFL defenses rely on similarity-based neighbor screening that requires every client to exchange and compare complete high-dimensional model vectors with all neighbors in each training round, creating prohibitive communication and computational costs that prevent deployment at web scale.
-We propose SketchGuard, a general framework that decouples Byzantine filtering from model aggregation through sketch-based neighbor screening.
-SketchGuard compresses $d$-dimensional models to $k$-dimensional sketches ($k\ll d$) using Count Sketch for similarity comparisons, then selectively fetches full models only from accepted neighbors, reducing per-round communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$, where $|\mathcal{N}_{i}|$ is the neighbor count and $|\mathcal{S}_{i}|\leq|\mathcal{N}_{i}|$ is the accepted neighbor count.
-We establish rigorous convergence guarantees in both strongly convex and non-convex settings, proving that Count Sketch compression preserves Byzantine resilience with controlled degradation bounds where approximation errors introduce only a $(1+O(\epsilon))$ factor in the effective threshold parameter.
-Comprehensive experiments across multiple datasets, network topologies, and attack scenarios demonstrate that SketchGuard maintains identical robustness to state-of-the-art methods, with mean test error rate deviation of only up to 0.35 percentage points, while reducing computation time by up to 82% and communication overhead by 50-70% depending on filtering effectiveness, with benefits scaling multiplicatively with model dimensionality and network connectivity.
-These results establish the viability of sketch-based compression as a fundamental enabler of robust DFL at web scale.
-Decentralized Learning, Byzantine Robustness, Dimensionality Reduction, Scalable Machine Learning
-copyright: rightsretainedjournalyear: 2026ccs: Security and privacy Systems security
-1. Introduction
+Abstract
+Decentralized Federated Learning enables privacy-preserving collaborative training without centralized servers but remains vulnerable to Byzantine attacks.
+Existing defenses require exchanging high-dimensional model vectors with all neighbors each round, creating prohibitive costs at scale.
+We propose SketchGuard, which decouples Byzantine filtering from aggregation via sketch-based screening.
+SketchGuard compresses $d$-dimensional models to $k$-dimensional sketches ($k\ll d$) using Count Sketch, then fetches full models only from accepted neighbors, reducing communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$, where $|\mathcal{N}_{i}|$ is the neighbor count and $|\mathcal{S}_{i}|\leq|\mathcal{N}_{i}|$ is the accepted count.
+We prove convergence in strongly convex and non-convex settings, showing that approximation errors introduce only a $(1+O(\epsilon))$ factor in the effective threshold.
+Experiments demonstrate SketchGuard maintains state-of-the-art robustness (mean TER deviation $\leq$0.5 percentage points) while reducing computation by up to 82% and communication by 50–70%.
+Machine Learning, ICML
+1 Introduction
 Federated Learning (FL) enables collaborative training of AI models over distributed data while preserving privacy by keeping raw data local (McMahan et al., 2017).
 However, the canonical, server-assisted architecture of FL centralizes aggregation of model parameters, creating a single point of failure, a communication bottleneck, and trust issues (Fang et al., 2024).
-These drawbacks have catalyzed Decentralized Federated Learning (DFL), where clients exchange model updates in a peer-to-peer manner over a dynamic communication graph, thereby improving scalability and resilience (Beltrán et al., 2023).
+These drawbacks have catalyzed Decentralized Federated Learning (DFL), where clients exchange model updates in a peer-to-peer manner over a dynamic communication graph, thereby improving scalability and resilience (Martínez Beltrán et al., 2023).
 A central challenge in DFL is byzantine robustness: the ability to withstand malicious clients that submit arbitrary or carefully crafted updates to poison training, induce consensus drift, or trigger targeted failures (Blanchard et al., 2017; Baruch et al., 2019).
 Unlike centralized FL where robust rules such as Krum (Blanchard et al., 2017), coordinatewise Median (Yin et al., 2018), or Trimmed-Mean (Wang et al., 2025) are applied once at a centralized server, DFL requires every client to aggregate its neighbors’ updates under local, graph-limited views, often with non-IID data and time-varying connectivity.
-To address this added complexity, most DFL defenses adopt local-consistency filters, where clients accept a neighbor’s update only if it is sufficiently similar to their own state, and then average over the accepted subset (Guo et al., 2021; He et al., 2022; Fang et al., 2024; Cajaraville-Aboy et al., 2024; El-Mhamdi et al., 2021).
+To address this added complexity, most DFL defenses adopt local-consistency filters, where clients accept a neighbor’s update only if it is sufficiently similar to their own state, and then average over the accepted subset (Guo et al., 2022; He et al., 2022; Fang et al., 2024; Cajaraville-Aboy et al., 2025; El-Mhamdi et al., 2021).
 These mechanisms provide convergence and robustness guarantees in both strongly convex and non-convex model training settings, yet suffer from a fundamental scalability bottleneck: clients must exchange and compare complete, high-dimensional model vectors with all neighbors in every round.
@@
   • Through comprehensive experiments across multiple datasets, network topologies, and attack scenarios, we demonstrate that our approach achieves identical robustness to state-of-the-art methods while reducing communication overhead by 50-70% and computation time by up to 82%.
-2. Preliminaries and Related Work
+2 Preliminaries and Related Work
 This section provides technical background on DFL protocols, Byzantine attack models, and compression techniques, and discusses related work in each area.
-2.1. DFL Problem Formulation and Protocol
+2.1 DFL Problem Formulation and Protocol
 Consider $n$ clients connected by an undirected graph $G=(V,E)$, where each client $i\in V$ possesses a private dataset $\mathcal{D}_{i}$ and maintains a local model $\mathbf{w}_{i}\in\mathbb{R}^{d}$.
 The collective objective is to minimize the average empirical loss:
-(1) $\min_{\mathbf{w}\in\mathbb{R}^{d}}F(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{w}),$
+$\min_{\mathbf{w}\in\mathbb{R}^{d}}F(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{w}),$ (1)
 where $f_{i}(\mathbf{w})=\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{i}}[\ell(\mathbf{w};\mathbf{x},y)]$ is the expected loss over client $i$’s data distribution.
 The DFL protocol alternates between two phases. First, clients perform local updates:
-(2) $\mathbf{w}_{i}^{t+1/2}=\mathbf{w}_{i}^{t}-\eta\nabla f_{i}(\mathbf{w}_{i}^{t}).$
+$\mathbf{w}_{i}^{t+1/2}=\mathbf{w}_{i}^{t}-\eta\nabla f_{i}(\mathbf{w}_{i}^{t}).$ (2)
 Then, each client aggregates neighbor models according to:
-(3) $\mathbf{w}_{i}^{t+1}=\alpha\mathbf{w}_{i}^{t+1/2}+(1-\alpha)\cdot\mathrm{AGG}_{i}\left(\{\mathbf{w}_{j}^{t+1/2}:j\in\mathcal{N}_{i}\}\right),$
+$\mathbf{w}_{i}^{t+1}=\alpha\mathbf{w}_{i}^{t+1/2}+(1-\alpha)\cdot\mathrm{AGG}_{i}\left(\{\mathbf{w}_{j}^{t+1/2}:j\in\mathcal{N}_{i}\}\right),$ (3)
 where $\mathcal{N}_{i}$ denotes client $i$’s neighbors and $\mathrm{AGG}_{i}$ is a local aggregation function.
-2.2. Byzantine Attack Models
+2.2 Byzantine Attack Models
 We consider $f$-Byzantine adversaries that control up to $f$ clients per neighborhood.
@@
 This constraint makes similarity-based filtering particularly attractive, as it provides a principled approach to neighbor screening that remains effective across different attack types while being computationally tractable for individual clients.
-2.3. Byzantine-Robust DFL Defenses
+2.3 Byzantine-Robust DFL Defenses
 As explained in the previous section, existing DFL defenses have converged on similarity-based neighbor filtering as the dominant paradigm, with methods differing primarily in their filtering criteria and aggregation mechanisms.
-UBAR (Guo et al., 2021) utilizes two-stage neighbor selection based on distance and loss filtering, followed by averaging.
+UBAR (Guo et al., 2022) utilizes two-stage neighbor selection based on distance and loss filtering, followed by averaging.
 LEARN (El-Mhamdi et al., 2021) employs multiple rounds of model exchanges per iteration with trimmed-mean aggregation.
@@
 It achieves state-of-the-art robustness but at a high computational cost.
-WFAgg (Cajaraville-Aboy et al., 2024) proposes multi-filter approaches for dynamic topologies, combining several screening mechanisms while maintaining the same computational complexity limitations.
+WFAgg (Cajaraville-Aboy et al., 2025) proposes multi-filter approaches for dynamic topologies, combining several screening mechanisms while maintaining the same computational complexity limitations.
 However, all these methods require computation of similarities between complete $d$-dimensional model vectors with all neighbors, creating bottlenecks that prevent practical web-scale deployments.
-Table 1 compares SketchGuard with these existing DFL defense approaches.
-Table 1.
-Comparison of Byzantine-robust decentralized aggregation algorithms.
-“Convex”/“Non-convex” indicate convergence guarantees; “No $c_{i}$” means no need to know compromised node ratio; “No Complete” means no complete-graph assumption.
-Algorithm Convex Non-convex No $c_{i}$ No Complete Scalable
-UBAR (Guo et al., 2021) Med.
-LEARN (El-Mhamdi et al., 2021) Low
-SCCLIP (He et al., 2022) Med.
-BALANCE (Fang et al., 2024) Med.
-WFAgg (Cajaraville-Aboy et al., 2024) Med.
-SketchGuard High
-2.4. Compression Techniques in FL
+A detailed comparison of SketchGuard with existing methods is provided in Appendix C.
+2.4 Compression Techniques in FL
 Communication and computational efficiency remain fundamental challenges in FL, motivating extensive research in model compression.
-Quantization techniques reduce parameter precision (Alistarh et al., 2017; Bernstein et al., 2018), sparsification methods transmit only significant updates (Stich et al., 2018; Sattler et al., 2019), and low-rank approaches exploit model structure (Konečnỳ et al., 2016; Haddadpour et al., 2021).
+Quantization techniques reduce parameter precision (Alistarh et al., 2017), sparsification methods transmit only significant updates (Stich et al., 2018; Sattler et al., 2020), and low-rank approaches exploit model structure (Konečný et al., 2016; Haddadpour et al., 2021).
 However, most compression techniques target benign settings and often compromise robustness when Byzantine participants are present.
@@
 More broadly, existing compression methods are incompatible with Byzantine-robust aggregation algorithms, preventing scalable deployment of robust FL.
-2.5. Count Sketch
+2.5 Count Sketch
 Count Sketch (Charikar et al., 2002) provides randomized linear projection with properties that make it particularly suitable for Byzantine-robust neighbor screening in DFL.
 Given $\mathbf{w}\in\mathbb{R}^{d}$, a Count Sketch of size $k\ll d$ is constructed using hash function $h:[d]\rightarrow[k]$ and sign function $s:[d]\rightarrow\{-1,+1\}$:
-(4) $\mathrm{CS}(\mathbf{w})[b]=\sum_{i:h(i)=b}s(i)w_{i},\quad b=1,\ldots,k.$
+$\mathrm{CS}(\mathbf{w})[b]=\sum_{i:h(i)=b}s(i)w_{i},\quad b=1,\ldots,k.$ (4)
 Count Sketch possesses three critical properties for robust DFL screening.
@@
 Third, and most crucial for similarity-based Byzantine filtering, Count Sketch preserves approximate Euclidean distances:
-Lemma 0 (Distance Preservation (Charikar et al., 2002)).
+Lemma 2.1 (Distance Preservation (Charikar et al., 2002)).
 For any vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{d}$ and sketch size $k=O(\epsilon^{-2}\log(1/\delta))$, with probability at least $1-\delta$:
-(5) $(1-\epsilon)\|\mathbf{u}-\mathbf{v}\|^{2}\leq\|\mathrm{CS}(\mathbf{u})-\mathrm{CS}(\mathbf{v})\|^{2}\leq(1+\epsilon)\|\mathbf{u}-\mathbf{v}\|^{2}$
-3. SketchGuard: Scalable Robust Aggregation
-We now present SketchGuard, which leverages Lemma 1 to perform neighbor screening in the sketch domain.
+$(1-\epsilon)\|\mathbf{u}-\mathbf{v}\|^{2}\leq\|\mathrm{CS}(\mathbf{u})-\mathrm{CS}(\mathbf{v})\|^{2}\leq(1+\epsilon)\|\mathbf{u}-\mathbf{v}\|^{2}$ (5)
+3 SketchGuard: Scalable Robust Aggregation
+We now present SketchGuard, which leverages Lemma 2.1 to perform neighbor screening in the sketch domain.
 The key design rationale is that if Count Sketch approximately preserves distances, then similarity-based filtering decisions made using sketches will closely match those made using full-precision models, allowing us to defer expensive full-model exchanges until after filtering.
-3.1. Protocol Description
-Figure 1. The SketchGuard Protocol:
+3.1 Protocol Description
+Figure 1: The SketchGuard Protocol
 At each DFL training round $t$, client $i$ executes the four-phase SketchGuard protocol detailed in Algorithm 1 and illustrated in Figure 1 as follows:
@@
 Phase 3: Adaptive Filtering. Neighbor $j$ is accepted if their sketch distance satisfies:
-(6) $\|\mathbf{s}_{i}^{t+1/2}-\mathbf{s}_{j}^{t+1/2}\|\leq\gamma\exp(-\kappa t/T)\|\mathbf{s}_{i}^{t+1/2}\|$
+$\|\mathbf{s}_{i}^{t+1/2}-\mathbf{s}_{j}^{t+1/2}\|\leq\gamma\exp(-\kappa t/T)\|\mathbf{s}_{i}^{t+1/2}\|$ (6)
 where $\gamma$ controls the base threshold, $\kappa$ the decay rate, and $T$ the total rounds.
 This adaptive threshold mechanism is adopted from BALANCE (Fang et al., 2024), where the exponential decay reflects convergence of honest clients over time.
-Remark 1.
-When using Count Sketch compression with approximation parameter $\epsilon$, the effective threshold parameter becomes $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$ to account for distance preservation errors in the sketch domain, as established by the analysis in Lemma 1.
+Remark 3.1.
+When using Count Sketch compression with approximation parameter $\epsilon$, the effective threshold parameter becomes $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$ to account for distance preservation errors in the sketch domain, as established by the analysis in Lemma 2.1.
 This ensures that our theoretical convergence guarantees in Section 4 accurately reflect the impact of compression on filtering decisions.
@@
 The verified models are then aggregated:
-(7) $\mathbf{w}_{i}^{t+1}=\alpha\mathbf{w}_{i}^{t+1/2}+\frac{(1-\alpha)}{|\mathcal{S}_{i}^{t}|}\sum_{j\in\mathcal{S}_{i}^{t}}\mathbf{w}_{j}^{t+1/2}$
+$\mathbf{w}_{i}^{t+1}=\alpha\mathbf{w}_{i}^{t+1/2}+\frac{(1-\alpha)}{|\mathcal{S}_{i}^{t}|}\sum_{j\in\mathcal{S}_{i}^{t}}\mathbf{w}_{j}^{t+1/2}$ (7)
 where $\alpha\in[0,1]$ balances self-reliance and collaboration.
@@
 12:  return $\mathbf{w}_{i}^{t+1}$
-3.2. Complexity Analysis and Performance Trade-offs
+3.2 Complexity Analysis and Performance Trade-offs
 We analyze the per-node-per-round complexity of SketchGuard compared to the best-case computation and communication complexity of existing Byzantine-robust methods.
@@
 This yields total complexity $O(d+k\cdot|\mathcal{N}_{i}|+d\cdot|\mathcal{S}_{i}^{t}|)$, compared to $O(d\cdot|\mathcal{N}_{i}|)$ for traditional methods that perform full-precision screening of all neighbors.
-Table 2. Computational complexity comparison with state-of-the-art (SOTA) across training phases.
+Table 1: Computational complexity comparison with state-of-the-art (SOTA) across training phases.
 Training Phase SOTA SketchGuard
@@
 Neighbor Screening $O(d\cdot|\mathcal{N}_{i}|)$ $O(k\cdot|\mathcal{N}_{i}|)$
-Model Verification & Aggregation $O(d\cdot|\mathcal{N}_{i}|)$ $O(d\cdot|\mathcal{S}_{i}^{t}|)$
-Total Per Round $O(d\cdot|\mathcal{N}_{i}|)$ $O(d+k\cdot|\mathcal{N}_{i}|+d\cdot|\mathcal{S}_{i}^{t}|)$
+Model Verification & Aggregation $O(d\cdot|\mathcal{S}_{i}^{t}|)$ $O(d\cdot|\mathcal{S}_{i}^{t}|)$
+Total Per Round $O(d+d\cdot|\mathcal{N}_{i}|+d\cdot|\mathcal{S}_{i}^{t}|)$ $O(d+k\cdot|\mathcal{N}_{i}|+d\cdot|\mathcal{S}_{i}^{t}|)$
 For communication complexity, SketchGuard transmits two types of messages: sketches to all neighbors and full models to accepted neighbors.
@@
 While sketch generation and exchange introduce $O(d+k\cdot|\mathcal{N}_{i}|)$ overhead even in benign settings, this cost becomes negligible as model dimensionality and network scale increase, which characterizes the web-scale deployments where Byzantine robustness is necessary.
-4. Convergence Analysis
+4 Convergence Analysis
 In this section, we establish that SketchGuard maintains convergence guarantees despite compressed filtering by proving that sketch-based screening preserves the robustness properties of full-precision similarity-based defenses with controlled degradation bounds.
-4.1. Technical Assumptions
+4.1 Technical Assumptions
 We state the standard assumptions that impact on convergence analysis.
-Assumption 1 (Strong Convexity).
+Assumption 4.1 (Strong Convexity).
 The population risk $F(\mathbf{w})$ is $\mu$-strongly convex, i.e., for all $\mathbf{w}_{1},\mathbf{w}_{2}\in\Theta$, one has that:
 $F(\mathbf{w}_{1})+\langle\nabla F(\mathbf{w}_{1}),\mathbf{w}_{2}-\mathbf{w}_{1}\rangle+\frac{\mu}{2}\|\mathbf{w}_{2}-\mathbf{w}_{1}\|^{2}\leq F(\mathbf{w}_{2}).$
-Assumption 2 (Smoothness).
+Assumption 4.2 (Smoothness).
 The population risk $F(\mathbf{w})$ is $L$-smooth, i.e., for all $\mathbf{w}_{1},\mathbf{w}_{2}\in\Theta$, we have that:
 $\|\nabla F(\mathbf{w}_{1})-\nabla F(\mathbf{w}_{2})\|\leq L\|\mathbf{w}_{1}-\mathbf{w}_{2}\|.$
-Assumption 3 (Bounded Variance).
+Assumption 4.3 (Bounded Variance).
 The stochastic gradient $\mathbf{g}(\mathbf{w}_{i})$ computed by an honest client $i\in\mathcal{H}$ is an unbiased estimator of the true gradient, and $\mathbf{g}(\mathbf{w}_{i})$ has bounded variance, where $\mathcal{H}$ is the set of honest clients.
@@
 $\mathbb{E}[\mathbf{g}(\mathbf{w}_{i})]=\nabla F(\mathbf{w}_{i}),\quad\mathbb{E}[\|\mathbf{g}(\mathbf{w}_{i})-\nabla F(\mathbf{w}_{i})\|^{2}]\leq\delta^{2}.$
-Assumption 4 (Bounded Parameters).
+Assumption 4.4 (Bounded Parameters).
 For any honest client $i\in\mathcal{H}$, the model $\mathbf{w}_{i}$ and $\|\nabla F(\mathbf{w}_{i})\|$ are bounded.
 That is, $\forall i\in\mathcal{H}$, we have $\|\mathbf{w}_{i}\|\leq\psi$, and $\|\nabla F(\mathbf{w}_{i})\|\leq\rho$.
-Assumption 5 (Graph Connectivity).
+Assumption 4.5 (Graph Connectivity).
 The subgraph induced by honest clients $G_{\mathcal{H}}$ remains connected throughout training.
-Assumption 6 (Hash Function Synchronization).
+Assumption 4.6 (Hash Function Synchronization).
 All clients use identical hash and sign functions for consistent sketching across the network.
-4.2. Main Convergence Results
-Leveraging Lemma 1, we establish convergence guarantees for SketchGuard that achieve similar results to those based on full-precision methods.
-Theorem 1 (Strongly Convex Convergence with Compression).
+This is trivially satisfied by seeding the hash functions with the model’s parameter count $d$, which is implicitly shared among all clients training the same architecture.
+4.2 Main Convergence Results
+Leveraging Lemma 2.1, we establish convergence guarantees for SketchGuard that achieve similar results to those based on full-precision methods.
+Theorem 4.7 (Strongly Convex Convergence with Compression).
 For $\mu$-strongly convex and $L$-smooth objectives, with learning rate $\eta\leq\min\{1/(4L),1/\mu\}$ and effective threshold parameter $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$, after $T$ rounds:
 $\mathbb{E}[F(\mathbf{w}_{i}^{T})-F(\mathbf{w}^{*})]\leq(1-\mu\eta)^{T}[F(\mathbf{w}_{i}^{0})-F(\mathbf{w}^{*})]$
-(8) $\quad+\frac{2L\eta\delta^{2}}{\mu}+\frac{2\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\mu\eta}$
+$\quad+\frac{2L\eta\delta^{2}}{\mu}+\frac{2\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\mu\eta}$ (8)
 where the compression error manifests only through the $(1+\epsilon)/(1-\epsilon)$ factor in $\gamma_{\text{eff}}$.
@@
 See Appendix A. ∎
-Theorem 2 (Non-Convex Convergence with Compression).
+Theorem 4.8 (Non-Convex Convergence with Compression).
 For non-convex $L$-smooth objectives, with the same parameter choices:
 $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}]\leq\frac{2[F(\mathbf{w}_{i}^{0})-F^{*}]}{\eta T}$
-(9) $\quad+4L\eta\delta^{2}+\frac{4\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\eta}$
+$\quad+4L\eta\delta^{2}+\frac{4\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\eta}$ (9)
 Proof.
@@
 For practical values of $\epsilon$ (e.g., $\epsilon=0.1$ yields $\gamma_{\text{eff}}\approx 1.1\gamma$), this degradation is minimal while enabling substantial compression.
-4.3. Robustness Preservation Under Compression
+4.3 Robustness Preservation Under Compression
 Our final theoretical result establishes that sketch compression preserves Byzantine robustness guarantees with controlled degradation.
-Lemma 0 (Sketch Compression Preserves Byzantine Resilience).
-Under the conditions of Theorems 1 and 2, sketch-based filtering with effective threshold $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$ provides equivalent Byzantine resilience to full-precision filtering with threshold $\gamma_{\text{eff}}$.
+Lemma 4.9 (Sketch Compression Preserves Byzantine Resilience).
+Under the conditions of Theorems 4.7 and 4.8, sketch-based filtering with effective threshold $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$ provides equivalent Byzantine resilience to full-precision filtering with threshold $\gamma_{\text{eff}}$.
 Specifically, compression does not create new attack vectors beyond the controlled threshold degradation.
 Proof.
-This follows directly from the convergence analysis in Theorems 1 and 2. Both proofs bound the neighbor difference terms under sketch-based filtering (Appendix A.5):
+This follows directly from the convergence analysis in Theorems 4.7 and 4.8. Both proofs bound the neighbor difference terms under sketch-based filtering (Appendix A.5):
 $\left\|\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|\leq\gamma_{\text{eff}}\psi$
@@
 ∎
-5. Performance Evaluation
-We evaluate SketchGuard [^1] Code and experiment artifacts available at doi.org/10.5281/zenodo.17223405.
+5 Performance Evaluation
+We evaluate SketchGuard [^1] Code and experiment artifacts available at [Anonymized Repository].
 through comprehensive experiments demonstrating that sketch-based compression maintains identical Byzantine robustness to state-of-the-art full-precision defenses while substantially reducing computational and communication overhead.
-5.1. Experimental Setup
-5.1.1. Datasets and Models
-We conduct experiments on two standard federated learning benchmarks from the LEAF suite (Caldas et al., 2018), chosen for their naturally non-IID data distributions that closely reflect real-world heterogeneity in federated environments.
-We evaluate both robustness and scalability across diverse model sizes.
-FEMNIST (Federated Extended MNIST) is a character recognition dataset with 62 classes distributed across 3,550 users, where each user represents a different writer’s handwriting samples.
-We employ a convolutional neural network comprising two convolutional layers followed by two fully connected layers.
-Our baseline configuration uses 6.6M parameters, which we systematically scale from 220K to over 60M parameters in our dimensionality scaling experiments (Section 5.3) to evaluate SketchGuard under varying model complexities representative of modern federated learning deployments.
-CelebA is a celebrity face attributes dataset distributed across 9,343 users.
-We use it for binary smile classification (smiling or not smiling), with a CNN architecture adapted for RGB inputs ($84\times 84$ pixels) containing 2.2M parameters.
-The network consists of two convolutional layers and two fully connected layers.
-Complete architectural specifications for both models are provided in Appendix C.1.
-5.1.2. Network Topologies
-(a) Ring(b) ER Sparsep=0.2p=0.2(c) ER Mediump=0.45p=0.45(d) ER Densep=0.6p=0.6(e) Fully Connected
-Figure 2. Network topologies used in the robustness evaluation experiments.
-We evaluate defense mechanisms across carefully selected network topologies that capture the spectrum of connectivity patterns in real-world decentralized systems, from resource-constrained sparse networks to well-connected infrastructures.
-For robustness evaluation experiments in Section 5.2, we employ Erdős-Rényi (ER) random graphs (Erdős and Rényi, 1960), which naturally model peer-to-peer networks where connections form probabilistically.
-ER graphs are generated by independently connecting each pair of nodes with probability $p$, producing networks with variable local connectivity that reflect real-world decentralized learning scenarios where participants discover neighbors organically rather than through engineered topologies.
-We evaluate five topologies spanning from minimal connectivity (Ring) to maximum connectivity (Fully Connected), with three intermediate ER graph densities as shown in Figure 2; complete specifications are provided in Appendix C.2.1.
-These experiments use 20 nodes to enable comprehensive evaluation across varying malicious client fractions and defense mechanisms while maintaining computational tractability.
-For network scaling experiments in Section 5.3, we employ k-regular graphs (Bondy et al., 1976) where each node has exactly $k$ neighbors.
-This design choice ensures consistent per-node computational load across all defense mechanisms, enabling fair comparison of algorithmic efficiency independent of topology-induced variance.
-We evaluate networks scaling from 20 to 300 nodes, demonstrating SketchGuard’s scalability to large-scale deployments with hundreds of participants.
-5.1.3. Attack Models
-We evaluate Byzantine robustness using two representative attack strategies that represent different threat models and attack sophistication levels.
-Directed Deviation attacks employ an optimization-based approach that crafts malicious updates by solving for parameters that maximize deviation in the direction opposite to honest gradient descent.
-This adaptive attack strategy represents a strong adversary that actively attempts to subvert Byzantine-robust aggregation mechanisms by carefully positioning malicious updates to evade detection while maximizing impact.
-We implement the directed deviation attack following the formulation in (Fang et al., 2020), which has demonstrated effectiveness against common Byzantine-robust aggregation rules including Krum and trimmed mean.
-Gaussian attacks (Blanchard et al., 2017) inject random noise $\hat{\mathbf{w}}_{j}=\mathbf{w}_{j}+\mathcal{N}(0,\sigma^{2}\mathbf{I})$ with $\sigma=1$, representing less sophisticated but realistic adversaries that disrupt convergence through stochastic perturbations.
-While simpler than directed deviation, these attacks model practical scenarios where attackers lack complete knowledge of the aggregation mechanism or resources for optimization-based attacks.
-We vary the fraction of Byzantine clients from 0% to 80% in 10% increments to evaluate robustness across different attack intensities.
-These attack strategies provide complementary evaluation: directed deviation establishes defense performance against sophisticated adaptive adversaries, while Gaussian attacks validate robustness against simpler but more common threat models.
-5.1.4. Baselines and Configuration
-We compare SketchGuard against four baseline methods: D-FedAvg (Lian et al., 2017), KRUM (Blanchard et al., 2017), UBAR (Guo et al., 2021), and BALANCE (Fang et al., 2024), representing the spectrum from non-robust to state-of-the-art Byzantine-robust aggregation.
-For all experiments, except the $k$-ablation study, we employ sketch sizes $k=1000$ for FEMNIST and $k=350$ for CelebA.
-For the baseline model configurations, this achieves compression ratios of 6603:1 and 6342:1 respectively, with approximation error $\epsilon\lesssim 0.2$ and probability $1-\delta>0.99$.
-All experiments use standard DFL configurations with 10 global rounds and 3 local epochs per round.
-We repeat each experiment across 3 random seeds and report mean results; observed variances are minimal across all metrics.
-Complete hyperparameter specifications are provided in Appendix C.2.2.
-5.1.5. Evaluation Metrics
-We evaluate defense mechanisms using three complementary metrics that capture robustness, computational efficiency, and communication overhead.
-Test Error Rate (TER). Defined as 1test accuracy1-$\text${test accuracy}, TER measures performance degradation averaged across honest clients. Lower values indicate better Byzantine resilience.
-Per-Round Computation Time.
-Wall-clock time for neighbor screening and aggregation phases per client per round, excluding local training which remains constant across methods.
-All measurements are recorded on identical hardware.
-Communication Overhead.
-Total parameters transmitted per client per round, including sketches and full model exchanges.
-We report both absolute counts and reduction percentages relative to full-precision baselines.
-5.2. Byzantine Robustness Evaluation
-5.2.1. Attack Strategy Resilience
-Figure 3. Impact of fraction of malicious clients on TER across different datasets and attack types.
-Figure 3 evaluates SketchGuard’s robustness against both directed deviation and Gaussian attacks on FEMNIST and CelebA datasets, averaged across all network topologies.
-SketchGuard achieves robustness equivalent to state-of-the-art methods BALANCE and UBAR across both attack strategies.
-Compared to BALANCE, SketchGuard exhibits statistically identical performance with mean absolute TER deviation of 0.02 percentage points and maximum deviation of 0.49 percentage points.
-Compared to UBAR, SketchGuard shows mean absolute TER deviation of 0.35 percentage points and maximum deviation of 1.51 percentage points, with SketchGuard having marginally higher TER on average by 0.20 percentage points; however, UBAR’s loss-based screening mechanism requires significantly higher computational overhead than both BALANCE and SketchGuard, as demonstrated in Section 5.3.
-Against both directed deviation (Figures 3a, 3c) and Gaussian attacks (Figures 3b, 3d), SketchGuard maintains consistently low error rates across all Byzantine client fractions, demonstrating that sketch-based compression preserves the robustness properties of full-precision similarity-based defenses while enabling substantial efficiency gains.
-5.2.2. Sketch Size Robustness
-SketchGuard demonstrates remarkable robustness to sketch size selection.
-Across $k$ values ranging from 500 to 100,000 on both FEMNIST and CelebA datasets under 50% Byzantine clients with directed deviation attacks, performance remains completely stable (TER = 15.59% for FEMNIST, 10.35% for CelebA across all $k$ values).
-Even with compression ratios exceeding 13,000:1 ($k=500$ for FEMNIST’s 6.6M parameters), SketchGuard preserves full Byzantine resilience.
-This insensitivity to $k$ stems from Count Sketch’s distance preservation guarantees: while smaller $k$ values introduce larger approximation errors, the resulting sketch-based distances remain sufficiently accurate to distinguish Byzantine node signatures from honest node behavior.
-The adaptive threshold mechanism $\gamma_{\text{eff}}$ compensates for increased approximation error, ensuring robust filtering across the entire practical range of sketch sizes.
-5.2.3. Topology-Dependent Robustness
-Figure 4. Impact of fraction of malicious clients on TER across different network topologies.
-Figure 4 presents SketchGuard’s robustness across five network topologies, averaged over attack types and datasets.
-SketchGuard demonstrates identical robustness to BALANCE and UBAR across all network topologies, from minimal connectivity in Ring topology to maximum connectivity in Fully Connected networks.
-In well-connected topologies (ER graphs with $p\geq 0.45$, Fully Connected), all three Byzantine-robust methods maintain low error rates below 20% even when 80% of the network is compromised, significantly outperforming KRUM and the non-robust D-FedAvg baseline.
-In extremely sparse topologies (Ring), BALANCE, UBAR, and SketchGuard exhibit graceful degradation, tolerating up to 40% Byzantine clients before substantial performance loss.
-Despite having only two neighbors per node in the Ring topology, SketchGuard maintains performance parity with full-precision methods across all attack intensities, demonstrating that sketch-based compression does not compromise filtering effectiveness even under severe connectivity constraints.
-5.3. Computational Efficiency
-Figure 5 evaluates SketchGuard’s computational efficiency through network size and model dimensionality scaling experiments on FEMNIST with 50% Byzantine clients under directed deviation attacks.
-Network scaling uses the baseline 6.6M parameter model across varying k-regular graph sizes, while model dimensionality scaling uses a fixed k-regular graph with node degree 154 across varying model sizes as detailed in Appendix C.3.
-Figure 5. Impact of network size and model dimensionality on per-node computation time.
-5.3.1. Network Size Scalability
-SketchGuard exhibits near-constant computational cost as network connectivity increases from 16 to 299 neighbors (Figure 5a), averaging 0.35s per round, while BALANCE and UBAR scale linearly with neighbor count.
-At smaller network sizes ($\leq$32 neighbors), SketchGuard’s sketch generation overhead dominates, resulting in higher absolute costs than full-precision methods.
-However, as connectivity increases beyond 96 neighbors, SketchGuard’s compressed screening operations yield substantial benefits.
-The efficiency gap widens multiplicatively with network size: at 154 neighbors, SketchGuard requires 0.34s compared to BALANCE’s 0.51s and UBAR’s 1.10s, achieving 33% and 69% savings respectively; at 299 neighbors, these savings expand to 60% and 82%, with SketchGuard requiring only 0.39s versus BALANCE’s 0.99s and UBAR’s 2.14s.
-This scalability advantage stems from neighbor screening operating in compressed $k$-dimensional space rather than full $d$-dimensional space, yielding multiplicative gains proportional to network connectivity.
-5.3.2. Model Dimensionality Scalability
-SketchGuard demonstrates sub-linear scaling as model size increases from 220K to 60M parameters (Figure 5b), contrasting sharply with the linear scaling of BALANCE and UBAR.
-For models exceeding 26M parameters, SketchGuard maintains average computational improvement of 64% over state-of-the-art methods.
-At 60M parameters, SketchGuard requires 2.0s versus BALANCE’s 4.2s and UBAR’s 10.0s, achieving 52% and 80% reductions respectively.
-This sub-linear scaling emerges from Count Sketch’s property that sketch size $k$ depends on approximation parameters ($\epsilon$, $\delta$) rather than model dimension $d$.
-While full-precision methods incur $O(d|\mathcal{N}_{i}|)$ comparison costs that grow linearly with model size, SketchGuard’s sketch comparison cost remains fixed at $O(k|\mathcal{N}_{i}|)$ regardless of $d$, proving particularly valuable for large-scale DFL scenarios involving billion-parameter models (Borzunov et al., 2023).
-5.4. Communication Efficiency
-SketchGuard reduces per-round communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$, where $|\mathcal{S}_{i}|\leq|\mathcal{N}_{i}|$ denotes accepted neighbors post-filtering.
-5.4.1. Compression Ratios and Bandwidth Reduction
-With compression ratios exceeding 6300:1 for both datasets, sketch transmission overhead becomes negligible compared to full model exchanges.
-Table 3 presents communication efficiency under different filtering scenarios.
-In scenarios where Byzantine filtering successfully rejects 70% of malicious neighbors, SketchGuard achieves approximately 70% communication reduction relative to full-precision methods.
-Even in benign settings where all neighbors are accepted, the sketch overhead of $O(k|\mathcal{N}_{i}|)$ represents less than 0.02% additional cost compared to the $O(d|\mathcal{N}_{i}|)$ baseline due to the compression ratio $k/d\approx 1/6300$, introducing negligible bandwidth penalty.
-Table 3. Communication Overhead Comparison
-Scenario Accepted Neighbors Reduction
-Benign (no filtering) $|\mathcal{S}_{i}|=|\mathcal{N}_{i}|$ <<0.02% penalty
-50% Byzantine filtering $|\mathcal{S}_{i}|\approx 0.5|\mathcal{N}_{i}|$ $\sim$50%
-70% Byzantine filtering $|\mathcal{S}_{i}|\approx 0.3|\mathcal{N}_{i}|$ $\sim$70%
-5.4.2. Multiplicative Scaling Benefits
-The communication savings scale multiplicatively with both model dimensionality and network connectivity.
-Consider a fully connected network with 100 neighbors per node and FEMNIST’s 6.6M parameter model under 50% Byzantine clients, where Count Sketch compresses models to 1K parameters:
-  • SketchGuard: Transmits 100K sketch parameters plus 330M full model parameters, totaling $\sim$330M parameters.
-  • Full-Precision Methods: Transmit 660M parameters to all neighbors.
-  • Reduction: $\sim$50% bandwidth savings.
-These bandwidth savings prove critical for resource-constrained devices and bandwidth-limited networks characteristic of real-world decentralized learning deployments.
-As model sizes grow and networks expand, SketchGuard’s communication efficiency becomes essential for practical deployment.
-6. Conclusions and Future Work
+5.1 Experimental Setup
+5.1.1 Datasets and Models
+We evaluate on three federated learning benchmarks from the LEAF suite (Caldas et al., 2019): FEMNIST (62-class character recognition, 3,550 users, CNN with 6.6M parameters), CelebA (binary smile classification, 9,343 users, CNN with 2.2M parameters), and Sent140 (Twitter sentiment analysis, 660,120 users, LSTM with 1.2M parameters).
+These datasets provide naturally non-IID distributions reflecting real-world federated heterogeneity across both image and text modalities.
+Complete architectural specifications are provided in Appendix D.1.
+5.1.2 Network Topologies
+We evaluate five topologies spanning from minimal connectivity (Ring, degree 2) to maximum connectivity (Fully Connected), with three intermediate Erdős-Rényi densities ($p\in\{0.2,0.45,0.6\}$).
+The Erdős-Rényi topologies are implemented as dynamic networks where connections are resampled after each round, modeling realistic peer-to-peer environments with changing connectivity.
+Robustness experiments use 20-node networks; scalability experiments use k-regular graphs scaling from 20 to 300 nodes, ensuring uniform node degree to isolate algorithmic scaling behavior from network topology effects.
+Complete specifications are in Appendix D.3.1.
+5.1.3 Attack Models and Evaluation Metrics
+We evaluate four attack strategies: Directed Deviation (Fang et al., 2020), an optimization-based attack maximizing deviation opposite to honest gradients; Gaussian (Blanchard et al., 2017), injecting random noise sampled from the Gaussian distribution; Krum (Fang et al., 2020), exploiting distance-based selection via clustered malicious models; and Backdoor (Bagdasaryan et al., 2020), embedding hidden trigger patterns for targeted misclassification.
+We vary Byzantine fraction from 0% to 80%.
+Detailed attack descriptions are in Appendix D.4.
+We report Test Error Rate (TER) for robustness, Attack Success Rate (ASR) for backdoor resilience, per-round computation time, and communication overhead.
+The details of these metrics are in Appendix D.2.
+5.1.4 Baselines and Configuration
+We compare against D-FedAvg (Lian et al., 2017), KRUM (Blanchard et al., 2017), UBAR (Guo et al., 2022), and BALANCE (Fang et al., 2024).
+We use sketch sizes $k=1000$ (FEMNIST), $k=350$ (CelebA), and $k=180$ (Sent140), maintaining compression ratios of approximately 6600:1 with $\epsilon\lesssim 0.2$.
+All experiments use 10 global rounds, 3 local epochs, and 3 random seeds.
+Complete hyperparameters are in Appendix D.3.2.
+5.2 Byzantine Robustness Evaluation
+Figure 2: Impact of fraction of malicious clients on TER across different datasets and attack types.
+SketchGuard maintains robustness equivalent to full-precision methods BALANCE and UBAR despite operating on compressed representations.
+Figure 2 evaluates whether sketch-based compression degrades Byzantine robustness compared to full-precision defenses.
+Across all four attack types on FEMNIST, CelebA, and Sent140 datasets, SketchGuard maintains robustness statistically indistinguishable from BALANCE and UBAR, with mean absolute TER deviation of only 0.38 percentage points versus BALANCE and 0.50 percentage points versus UBAR.
+This confirms that Count Sketch’s distance preservation properties (Lemma 2.1) translate to equivalent filtering decisions in practice.
+Against directed deviation and Krum attacks, all three similarity-based methods (BALANCE, UBAR, SketchGuard) maintain TER below 20% on FEMNIST and below 13% on CelebA even at 80% Byzantine clients, while D-FedAvg and KRUM collapse to near-random accuracy (TER >> 60%).
+The Krum defense fails against its targeted attack (TER >> 60% at 30%+ Byzantine), whereas similarity-based filtering—whether on full models or sketches—remains effective.
+For Gaussian attacks, SketchGuard matches BALANCE within 0.03 percentage points on average.
+For backdoor attacks, SketchGuard achieves ASR comparable to BALANCE and UBAR across all datasets.
+On FEMNIST, all three methods achieve low ASR (6.98–9.94%), demonstrating that similarity-based filtering effectively identifies the scaled model updates characteristic of backdoor attacks.
+On CelebA and Sent140, all robust methods substantially outperform D-FedAvg.
+These results confirm that sketch compression does not create vulnerabilities to backdoor attacks.
+5.3 Computational Efficiency
+Figure 3: Impact of network size and model dimensionality on per-node computation time. SketchGuard achieves substantial savings while maintaining equivalent robustness (Section 5.2).
+Having established that SketchGuard matches state-of-the-art robustness, we now demonstrate the efficiency gains that motivate sketch-based filtering.
+Figure 3 evaluates computational scaling on FEMNIST with 50% Byzantine clients under directed deviation attacks.
+SketchGuard exhibits near-constant cost as network connectivity increases from 16 to 299 neighbors (Figure 3a), averaging 0.35s per round, while BALANCE and UBAR scale linearly.
+At 299 neighbors, SketchGuard achieves 60% savings versus BALANCE (0.39s vs 0.99s) and 82% versus UBAR (0.39s vs 2.14s).
+This advantage grows with network density because sketch comparisons cost $O(k)$ per neighbor versus $O(d)$ for full-precision methods.
+For model dimensionality scaling (Figure 3b), SketchGuard demonstrates sub-linear growth from 220K to 60M parameters, while BALANCE and UBAR scale linearly.
+At 60M parameters, SketchGuard requires 2.0s versus 4.2s (BALANCE) and 10.0s (UBAR)—52% and 80% reductions.
+This sub-linear scaling emerges because sketch size $k$ depends on approximation parameters ($\epsilon$, $\delta$) rather than model dimension $d$, making SketchGuard increasingly valuable as models grow toward billion-parameter scales (Borzunov et al., 2023).
+5.4 Communication Efficiency
+SketchGuard reduces per-round communication from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$, where $|\mathcal{S}_{i}|\leq|\mathcal{N}_{i}|$ is the accepted neighbor count.
+With compression ratios of approximately 6600:1, sketch overhead is negligible (<<0.02% in benign settings).
+When Byzantine filtering rejects 50–70% of neighbors, communication reductions of 50–70% follow directly.
+These savings scale multiplicatively with model size and network connectivity, addressing the fundamental scalability bottleneck that has limited practical deployment of Byzantine-robust DFL.
+5.5 Sensitivity Analysis
+5.5.1 Sketch Size Robustness
+A practical concern for sketch-based methods is sensitivity to the sketch size parameter $k$.
+SketchGuard demonstrates remarkable insensitivity: across $k$ values ranging from 500 to 100,000 on both FEMNIST and CelebA under 50% Byzantine clients with directed deviation attacks, TER remains completely stable (15.59% for FEMNIST, 10.35% for CelebA).
+Even at compression ratios exceeding 13,000:1 ($k=500$ for FEMNIST’s 6.6M parameters), robustness is fully preserved.
+This insensitivity stems from Count Sketch’s distance preservation guarantees: sketch-based distances remain sufficiently accurate to distinguish Byzantine updates from honest behavior across a wide range of compression levels.
+5.5.2 Topology Robustness
+Table 2 presents robustness across five network topologies for the three similarity-based methods (non-robust D-FedAvg and KRUM consistently fail at $\sim$63% TER as shown in Figure 2).
+SketchGuard matches BALANCE across all topologies, confirming that sketch compression does not interact negatively with network structure.
+In well-connected topologies (ER $p\geq 0.45$, Fully Connected), all methods maintain TER below 20% even at 80% Byzantine.
+In sparse Ring topology, all methods exhibit graceful degradation at 60%+ attack intensity due to limited honest neighbors—a fundamental limitation of local aggregation rather than any method-specific weakness.
+Table 2: Test Error Rate (%) across topologies for robust methods, averaged over datasets and attacks (excl. backdoor).
+Topology Byz.% UBAR BALANCE SG
+Ring 20 19.0 19.9 19.9
+40 18.9 19.8 19.6
+60 30.2 30.6 30.3
+80 64.3 63.7 64.2
+ER (pp=0.2) 20 20.6 20.8 20.7
+40 20.2 20.4 20.3
+60 19.4 19.8 19.7
+80 18.8 20.4 20.0
+ER (pp=0.45) 20 20.1 20.4 20.3
+40 20.4 20.1 19.9
+60 19.6 19.8 19.5
+80 19.2 20.3 20.2
+ER (pp=0.6) 20 20.1 20.6 20.3
+40 20.1 20.2 20.1
+60 19.7 19.7 19.5
+80 19.1 19.2 19.0
+Fully Conn. 20 20.0 20.5 20.3
+40 19.9 20.1 19.9
+60 19.5 19.7 19.5
+80 19.1 19.3 18.8
+6 Conclusions
 We proposed SketchGuard, a framework for scaling Byzantine-robust decentralized federated learning through sketch-based neighbor screening.
-By decoupling Byzantine filtering from model aggregation, our approach reduces communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$ while preserving robustness guarantees of full-precision similarity-based defenses.
-Our theoretical analysis establishes that Count Sketch compression maintains Byzantine resilience with controlled degradation bounds, where approximation errors introduce only a $(1+O(\epsilon))$ factor in the effective threshold parameter.
-We prove convergence guarantees in both strongly convex and non-convex settings, achieving optimal rates for their respective problem classes.
-Comprehensive experiments demonstrate that SketchGuard maintains identical robustness to state-of-the-art methods, with mean TER deviation of only up to 0.35 percentage points.
-This robustness equivalence is achieved while reducing computation time by up to 82%, with efficiency gains scaling multiplicatively with model dimensionality and network connectivity.
-Communication overhead scales proportionally with the fraction of accepted neighbors, achieving reductions of 50-70% under typical Byzantine attack scenarios.
-SketchGuard addresses the fundamental scalability bottleneck limiting real-world deployment of robust decentralized learning.
-Future work includes efficient hash synchronization protocols for dynamic networks, adaptive sketch sizing based on attack intensity, and theoretical extensions to non-IID data distributions.
-As FL expands into web-scale deployments with untrusted participants, scalable robustness mechanisms like SketchGuard become essential infrastructure for maintaining both security and performance in distributed machine learning systems.
-Acknowledgements.
-This work is supported by the Australian Research Council (ARC) through Discovery Project grant DP240102088 and Linkage Infrastructure, Equipment and Facilities grant LE200100049.
-Appendix A Proof of Theorem 1
+By decoupling filtering from aggregation, our approach reduces communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$ while preserving the robustness guarantees of full-precision defenses.
+Our theoretical analysis establishes that Count Sketch compression maintains Byzantine resilience with controlled degradation bounds, introducing only a $(1+O(\epsilon))$ factor in the effective threshold.
+Experiments confirm that SketchGuard matches state-of-the-art robustness (mean TER deviation $\leq$0.5 percentage points) while reducing computation by up to 82% and communication by 50–70% under typical attack scenarios.
+Impact Statement
+This paper presents work whose goal is to advance the field of Distributed Machine Learning.
+There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix A Proof of Theorem 4.7
 We provide a detailed proof of the convergence guarantee for SketchGuard under strongly convex objectives with sketch-based compression.
 The proof structure follows the analysis framework of BALANCE (Fang et al., 2024), with key modifications to handle Count Sketch compression errors through the effective threshold parameter $\gamma_{\text{eff}}$.
-A.1. Setup and Notation
+A.1 Setup and Notation
 We denote by $\mathbf{g}(\mathbf{w}_{i}^{t})$ the stochastic gradient computed by client $i$ at round $t$. The local update is:
-(10) $\mathbf{w}_{i}^{t+1/2}=\mathbf{w}_{i}^{t}-\eta\mathbf{g}(\mathbf{w}_{i}^{t}),$
+$\mathbf{w}_{i}^{t+1/2}=\mathbf{w}_{i}^{t}-\eta\mathbf{g}(\mathbf{w}_{i}^{t}),$ (10)
 where $\eta>0$ is the learning rate. For simplicity in the proof, we drop the superscript $t$ in $\mathcal{S}_{i}^{t}$ and write $\mathcal{S}_{i}$.
-A.2. Model Update Analysis
+A.2 Model Update Analysis
 The model update for client $i$ can be expressed as:
 $\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}$ $=\alpha\mathbf{w}_{i}^{t+1/2}$
-(11) $\quad+(1-\alpha)\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t}$
+$\quad+(1-\alpha)\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t}$ (11)
 $=\alpha\mathbf{w}_{i}^{t+1/2}+(1-\alpha)\frac{1}{|\mathcal{S}_{i}|}$
-(12) $\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2}+\mathbf{w}_{i}^{t+1/2})-\mathbf{w}_{i}^{t}$
+$\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2}+\mathbf{w}_{i}^{t+1/2})-\mathbf{w}_{i}^{t}$ (12)
 $=\mathbf{w}_{i}^{t+1/2}+\frac{1-\alpha}{|\mathcal{S}_{i}|}$
-(13) $\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})-\mathbf{w}_{i}^{t}$
+$\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})-\mathbf{w}_{i}^{t}$ (13)
 $=-\eta\mathbf{g}(\mathbf{w}_{i}^{t})+\frac{1-\alpha}{|\mathcal{S}_{i}|}$
-(14) $\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})$
-A.3. Applying Smoothness
+$\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})$ (14)
+A.3 Applying Smoothness
 By the $L$-smoothness assumption, we have:
 $F(\mathbf{w}_{i}^{t+1})$ $\leq F(\mathbf{w}_{i}^{t})+\langle\nabla F(\mathbf{w}_{i}^{t}),\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\rangle$
-(15) $\quad+\frac{L}{2}\|\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\|^{2}$
+$\quad+\frac{L}{2}\|\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\|^{2}$ (15)
 Substituting the model update expression:
@@
 $\quad+\frac{L}{2}\left\|-\eta\mathbf{g}(\mathbf{w}_{i}^{t})\right$.
-(16) $\quad\left.+\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$
-A.4. Bounding the Quadratic Term
+$\quad\left.+\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$ (16)
+A.4 Bounding the Quadratic Term
 Using the inequality $\|a+b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$:
@@
 $\leq 2\eta^{2}\|\mathbf{g}(\mathbf{w}_{i}^{t})\|^{2}$
-(17) $\quad+2\left\|\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$
-A.5. Impact of Sketch-Based Filtering
+$\quad+2\left\|\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$ (17)
+A.5 Impact of Sketch-Based Filtering
 With sketch-based filtering, neighbor $j$ is accepted if:
 $\|\mathrm{CS}(\mathbf{w}_{i}^{t+1/2})-\mathrm{CS}(\mathbf{w}_{j}^{t+1/2})\|$
-(18) $\quad\leq\gamma\exp(-\kappa t/T)\|\mathrm{CS}(\mathbf{w}_{i}^{t+1/2})\|$
+$\quad\leq\gamma\exp(-\kappa t/T)\|\mathrm{CS}(\mathbf{w}_{i}^{t+1/2})\|$ (18)
 By the distance preservation property of Count Sketch (Charikar et al., 2002), this implies:
 $\|\mathbf{w}_{i}^{t+1/2}-\mathbf{w}_{j}^{t+1/2}\|$
-(19) $\quad\leq\gamma\sqrt{\frac{1+\epsilon}{1-\epsilon}}\exp(-\kappa t/T)\|\mathbf{w}_{i}^{t+1/2}\|$
+$\quad\leq\gamma\sqrt{\frac{1+\epsilon}{1-\epsilon}}\exp(-\kappa t/T)\|\mathbf{w}_{i}^{t+1/2}\|$ (19)
 Define $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$. Then:
-(20) $\left\|\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|\leq\gamma_{\text{eff}}\psi$
+$\left\|\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|\leq\gamma_{\text{eff}}\psi$ (20)
 where we used the bounded parameters assumption that $\|\mathbf{w}_{i}^{t+1/2}\|\leq\psi$.
-A.6. Taking Expectation
+A.6 Taking Expectation
 Taking expectation and using the bounded variance assumption:
@@
 $\quad+L\eta^{2}(\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}+\delta^{2})$
-(21) $\quad+L(1-\alpha)^{2}\gamma_{\text{eff}}^{2}\psi^{2}$
-A.7. Simplifying with Parameter Constraints
+$\quad+L(1-\alpha)^{2}\gamma_{\text{eff}}^{2}\psi^{2}$ (21)
+A.7 Simplifying with Parameter Constraints
 With $\eta\leq 1/(4L)$, we have $L\eta^{2}\leq\eta/4$:
 $\mathbb{E}[F(\mathbf{w}_{i}^{t+1})]$ $\leq\mathbb{E}[F(\mathbf{w}_{i}^{t})]-\frac{\eta}{2}\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}$
-(22) $\quad+L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
+$\quad+L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$ (22)
 where we chose $\gamma_{\text{eff}}\leq\rho/(L\psi(1-\alpha))$.
-A.8. Applying Strong Convexity
+A.8 Applying Strong Convexity
 By the strong convexity assumption and the Polyak-Łojasiewicz inequality:
-(23) $\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}\geq 2\mu(F(\mathbf{w}_{i}^{t})-F(\mathbf{w}^{*}))$
+$\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}\geq 2\mu(F(\mathbf{w}_{i}^{t})-F(\mathbf{w}^{*}))$ (23)
 Therefore:
@@
 $\quad\leq(1-\mu\eta)\mathbb{E}[F(\mathbf{w}_{i}^{t})-F(\mathbf{w}^{*})]$
-(24) $\quad+L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
-A.9. Telescoping and Final Bound
+$\quad+L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$ (24)
+A.9 Telescoping and Final Bound
 Telescoping over $t=0,1,\ldots,T-1$:
@@
 $\quad\leq(1-\mu\eta)^{T}[F(\mathbf{w}_{i}^{0})-F(\mathbf{w}^{*})]$
-(25) $\quad+\frac{2L\eta\delta^{2}}{\mu}+\frac{2\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\mu\eta}$
-This completes the proof of Theorem 1.
+$\quad+\frac{2L\eta\delta^{2}}{\mu}+\frac{2\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\mu\eta}$ (25)
+This completes the proof of Theorem 4.7.
 The key observation is that the compression error only affects the convergence through $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$, introducing a controllable degradation factor while maintaining the same convergence rate as the state-of-the-art.
-Appendix B Proof of Theorem 2
+Appendix B Proof of Theorem 4.8
 We establish the convergence guarantee for SketchGuard in non-convex settings, adapting the BALANCE analysis framework (Fang et al., 2024) for sketch compression.
-B.1. Starting from Smoothness
+B.1 Starting from Smoothness
 Following the analysis from Appendix A up to the expectation bound, we have:
 $\mathbb{E}[F(\mathbf{w}_{i}^{t+1})]$ $\leq\mathbb{E}[F(\mathbf{w}_{i}^{t})]-\frac{\eta}{2}\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}$
-(26) $\quad+2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
-B.2. Rearranging for Gradient Norm
+$\quad+2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$ (26)
+B.2 Rearranging for Gradient Norm
 Rearranging the inequality:
@@
 $\quad\leq\mathbb{E}[F(\mathbf{w}_{i}^{t})-F(\mathbf{w}_{i}^{t+1})]$
-(27) $\quad+2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
-B.3. Telescoping
+$\quad+2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$ (27)
+B.3 Telescoping
 Summing from $t=0$ to T1T-1:
@@
 $\quad\leq F(\mathbf{w}_{i}^{0})-\mathbb{E}[F(\mathbf{w}_{i}^{T})]$
-(28) $\quad+T(2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha))$
-B.4. Averaging and Using Lower Bound
+$\quad+T(2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha))$ (28)
+B.4 Averaging and Using Lower Bound
 Since $F(\mathbf{w}_{i}^{T})\geq F^{*}$:
@@
 $\quad\leq\frac{2(F(\mathbf{w}_{i}^{0})-F^{*})}{\eta T}$
-(29) $\quad+4L\eta\delta^{2}+\frac{4\gamma_{\text{eff}}\psi\rho(1-\alpha)}{\eta}$
-This establishes Theorem 2. The convergence rate matches the optimal rate for non-convex optimization, with the compression error appearing only through $\gamma_{\text{eff}}$.
-Appendix C Supplementary Experimental Details
-This appendix provides comprehensive details about the experimental setup, including model architectures, hyperparameter configurations, network topologies, and scalability experiment specifications used throughout our study.
-C.1. Detailed Model Architectures
-We provide complete architectural specifications for the two datasets used in our experiments: FEMNIST and CelebA.
-Both architectures follow standard designs adapted to their respective input dimensions and task requirements.
-C.1.1. FEMNIST Architecture
+$\quad+4L\eta\delta^{2}+\frac{4\gamma_{\text{eff}}\psi\rho(1-\alpha)}{\eta}$ (29)
+This establishes Theorem 4.8. The convergence rate matches the optimal rate for non-convex optimization, with the compression error appearing only through $\gamma_{\text{eff}}$.
+Appendix C Comparison with Existing Methods
+Table 3 compares SketchGuard with existing Byzantine-robust DFL algorithms across key properties.
+SketchGuard is the only method that combines convergence guarantees in both convex and non-convex settings, operates without requiring knowledge of the compromised node ratio or complete graph assumptions, and achieves high scalability through sketch-based compression.
+Table 3: Comparison of Byzantine-robust decentralized aggregation algorithms.
+”Convex”/”Non-convex” indicate convergence guarantees; ”No $c_{i}$” means no need to know compromised node ratio; ”No Complete” means no complete-graph assumption.
+Algorithm Convex Non-convex No $c_{i}$ No Complete Scalable
+UBAR (Guo et al., 2022) Med.
+LEARN (El-Mhamdi et al., 2021) Low
+SCCLIP (He et al., 2022) Med.
+BALANCE (Fang et al., 2024) Med.
+WFAgg (Cajaraville-Aboy et al., 2025) Med.
+SketchGuard High
+Appendix D Supplementary Experimental Details
+This appendix provides comprehensive details about the experimental setup, including model architectures, hyperparameter configurations, network topologies, attack configurations, and scalability experiment specifications used throughout our study.
+D.1 Detailed Model Architectures
+We provide complete architectural specifications for the three datasets used in our experiments: FEMNIST, CelebA, and Sent140.
+The image-based models (FEMNIST and CelebA) follow standard convolutional designs, while Sent140 uses an LSTM-based architecture for sequential text data.
+D.1.1 FEMNIST Architecture
 The FEMNIST model follows the original LEAF specification with a convolutional neural network design.
@@
 Table 4 provides the complete layer-by-layer specification, including parameter counts and output dimensions.
-Table 4. FEMNIST Model Architecture Details
+Table 4: FEMNIST Model Architecture Details
 Layer Type Parameters Output Size
@@
 Total 6,603,710
-C.1.2. CelebA Architecture
+D.1.2 CelebA Architecture
 The CelebA model uses a LeNet-style architecture adapted for larger RGB input images ($84\times 84$ pixels).
@@
 The complete architecture is detailed in Table 5.
-Table 5. CelebA Model Architecture Details
+Table 5: CelebA Model Architecture Details
 Layer Type Parameters Output Size
@@
 Total 2,219,692
-C.2. Complete Experimental Configuration
+D.1.3 Sent140 Architecture
+The Sent140 model uses a stacked LSTM architecture following the LEAF reference implementation.
+The model processes variable-length text sequences through an embedding layer, a 2-layer LSTM, and fully connected output layers.
+Table 6 provides the complete specification.
+Table 6: Sent140 Model Architecture Details
+Layer Type Configuration Output Size
+Input seq_len
+Embedding Embedding vocab=10000, dim=100 $\text{seq\_len}\times 100$
+LSTM Stacked LSTM 2 layers, hidden=100 $100$
+Dropout Dropout $p=0.1$ $100$
+FC1 Linear $100\rightarrow 128$ $128$
+FC2 Linear $128\rightarrow 2$ 22
+Total $\sim$1.2M
+The embedding layer maps vocabulary indices (size 10,000) to 100-dimensional dense vectors.
+The stacked 2-layer LSTM with 100 hidden units processes the embedded sequence with dropout for regularization.
+The final hidden state passes through two fully connected layers for binary sentiment classification.
+D.2 Evaluation Metrics
+We evaluate defense mechanisms using four complementary metrics that capture robustness, backdoor resilience, computational efficiency, and communication overhead.
+Test Error Rate (TER).
+Defined as 1test accuracy1-$\text${test accuracy}, TER measures performance degradation averaged across honest clients.
+Lower values indicate better Byzantine resilience against model poisoning attacks.
+TER is computed on each honest client’s local test set and averaged across all honest clients in the network.
+Attack Success Rate (ASR).
+For backdoor attacks, ASR measures the fraction of triggered test samples classified as the attacker’s target label on honest client models.
+Lower ASR indicates better resilience to backdoor injection.
+The random-chance baseline is ASR $\approx 1/C$ where $C$ is the number of classes: 1.6% for FEMNIST (62 classes) and 50% for CelebA and Sent140 (binary classification).
+ASR is computed by applying the backdoor trigger to clean test samples and measuring the fraction misclassified to the target label.
+Per-Round Computation Time.
+Wall-clock time for neighbor screening and aggregation phases per client per round, excluding local training which remains constant across methods.
+This metric captures the computational overhead introduced by Byzantine-robust aggregation mechanisms.
+All timing measurements are recorded on identical hardware (CPU-only for reproducibility) and averaged across all clients and rounds.
+Communication Overhead.
+Total parameters transmitted per client per round, including both sketches (for SketchGuard) and full model exchanges.
+We report both absolute parameter counts and reduction percentages relative to full-precision baselines that transmit complete models to all neighbors.
+D.3 Complete Experimental Configuration
 This section details all hyperparameters and configuration settings used across our experiments to ensure reproducibility.
-C.2.1. Network Topology Specifications
+D.3.1 Network Topology Specifications
 We evaluate defense mechanisms across five different network topologies representing varying levels of connectivity.
@@
 Table 7 specifies the parameters and expected node degrees for each topology in our 20-node experimental networks.
-Table 6. Network Topology Configurations
+Table 7: Network Topology Configurations
 Topology Parameters Expected Degree
@@
 Fully Connected 19
-C.2.2. Hyperparameter Settings
+D.3.2 Hyperparameter Settings
 Table 8 presents the complete set of hyperparameters used in our experiments, organized by category.
@@
 Defense mechanism parameters ($\gamma$, $\kappa$, $\alpha$) were tuned to balance robustness and model utility.
-For SketchGuard, we set the sketch size $k=1000$ for FEMNIST and $k=350$ for CelebA to get roughly the same factor of compression for both datasets.
-Table 7. Complete Hyperparameter Configuration
+For SketchGuard, we set dataset-specific sketch sizes to achieve comparable compression ratios across different model dimensions.
+Table 8: Complete Hyperparameter Configuration
 Parameter Value
@@
 Number of clients 20
-Global epochs 10
+Global rounds 10
 Local epochs per round 3
 Batch size 64
-Learning rate 0.01 (FEMNIST), 0.001 (CelebA)
-Maximum samples per client 10,000 (FEMNIST), 4,500 (CelebA)
-Random seed 987654321, 39573295, 32599368
+Learning rate 0.01 (FEMNIST, Sent140), 0.001 (CelebA)
+Maximum samples per client 10,000 (FEMNIST), 4,500 (CelebA, Sent140)
+Random seeds 987654321, 39573295, 32599368
 SketchGuard Parameters
-Sketch size ($k$) 1,000 (FEMNIST), 350 (CelebA)
+Sketch size ($k$) 1,000 (FEMNIST), 350 (CelebA), 180 (Sent140)
 Threshold parameter ($\gamma$) 2.0
@@
 Compromised fraction Attack percentage
-C.3. Scalability Experiments
+D.3.3 Backdoor Attack Configuration
+Table 9 details the backdoor attack configurations used for each dataset.
+For image datasets (FEMNIST and CelebA), the backdoor trigger consists of a white square pattern inserted in the bottom-right corner of the image.
+For Sent140, the trigger consists of rare vocabulary token indices appended to the end of the text sequence.
+Table 9: Backdoor Attack Configuration by Dataset
+Dataset Target Label Trigger Size Trigger Type
+FEMNIST 0 $4\times 4$ pixels White square (bottom-right)
+CelebA 0 (not smiling) $8\times 8$ pixels White square (bottom-right)
+Sent140 1 (positive) 3 tokens Rare word indices appended
+For Sent140, the trigger tokens are selected from high vocabulary indices (9997, 9998, 9999) that rarely appear in natural text, ensuring the trigger pattern is distinctive and unlikely to occur in benign samples.
+D.4 Detailed Attack Descriptions
+We evaluate Byzantine robustness using four representative attack strategies that span different threat models and sophistication levels.
+D.4.1 Directed Deviation Attack
+Directed deviation attacks (Fang et al., 2020) employ an optimization-based approach that crafts malicious updates by solving for parameters that maximize deviation in the direction opposite to honest gradient descent.
+This adaptive attack strategy represents a strong adversary that actively attempts to subvert Byzantine-robust aggregation mechanisms by carefully positioning malicious updates to evade detection while maximizing impact.
+The attack computes the mean of honest updates and crafts malicious updates as:
+$\hat{\mathbf{w}}_{j}=\bar{\mathbf{w}}-\lambda\cdot\text{sign}(\bar{\mathbf{w}}-\mathbf{w}^{t})$ (30)
+where $\bar{\mathbf{w}}$ is the mean honest update, $\mathbf{w}^{t}$ is the current model, and $\lambda$ controls attack intensity.
+This attack has demonstrated effectiveness against common Byzantine-robust aggregation rules including Krum and trimmed mean.
+D.4.2 Gaussian Attack
+Gaussian attacks (Blanchard et al., 2017) inject random noise $\hat{\mathbf{w}}_{j}=\mathcal{N}(0,200\mathbf{I})$, representing less sophisticated but realistic adversaries that disrupt convergence through stochastic perturbations.
+While simpler than directed deviation, these attacks model practical scenarios where attackers lack complete knowledge of the aggregation mechanism or resources for optimization-based attacks.
+The high variance ($\sigma^{2}=200$) ensures significant disruption to model convergence.
+D.4.3 Krum Attack
+Krum attacks (Fang et al., 2020) craft malicious models that cause Krum to select the attacker’s chosen model. The attack works by:
+  1. 1.
+    Computing the mean of honest models as a reference
+  2. 2.
+    Creating a primary malicious model in the opposite direction of honest gradient descent
+  3. 3.
+    Creating supporting malicious models very close to the primary to minimize the Krum score of the primary model
+The attack exploits Krum’s reliance on pairwise distances: if attackers control enough nodes, they can cluster their models tightly while being far from honest models, causing Krum to select the malicious model as it appears most “central.”
+D.4.4 Backdoor Attack
+Backdoor attacks (Bagdasaryan et al., 2020) represent a distinct threat model where adversaries aim to embed hidden functionality rather than degrade overall performance.
+Malicious clients augment their training data with trigger patterns mapped to an attacker-chosen target label, then scale their model updates by a factor proportional to the total number of clients to amplify the backdoor’s persistence during aggregation.
+Unlike other attacks that maximize test error, backdoor attacks preserve main task accuracy while causing misclassification only on triggered inputs, making them particularly stealthy.
+Dataset-specific trigger configurations are detailed in Table 9.
+D.5 Scalability Experiments
 We conducted two types of scalability experiments to evaluate how defense mechanisms perform under different scaling conditions: network size scaling and model dimensionality scaling.
-C.3.1. Network Size Scaling
+D.5.1 Network Size Scaling
 To evaluate the computational scalability of defense mechanisms with respect to network size, we conducted experiments on k-regular graphs with varying numbers of participants.
@@
 These experiments used shorter training runs (3 rounds with 1 local epoch each) to focus on measuring computational overhead rather than convergence behavior.
-Table 8. Network Size Scaling Configurations
-Node Degree Network Size Attack Percentage
+All timing measurements were performed on CPU to ensure consistent and reproducible results across different hardware configurations.
+Table 10: Network Size Scaling Configurations
+Node Degree ($k$) Network Size Attack Percentage
 16 20 50%
@@
 299 300 50%
-C.3.2. Model Scaling Variants
+D.5.2 Model Scaling Variants
 To investigate the impact of model dimensionality on defense mechanism performance, we created five variants of the FEMNIST architecture with different parameter counts.
@@
 These variants were created by systematically scaling the number of convolutional filters and fully connected layer dimensions while maintaining the overall architectural structure.
-Table 9. Model Scaling Variants
+Table 11: Model Scaling Variants
 Variant Architecture Modifications Parameters
@@
 Small Standard configuration 848,382
-Standard Baseline FEMNIST architecture 6,603,710
+Baseline Baseline FEMNIST architecture 6,603,710
 Large Increased filter counts and hidden units 26,154,814
 XLarge Further increased dimensions 60,271,678
+Model dimensionality scaling experiments use a fixed k-regular graph with node degree 154 (155 total nodes) to isolate the effect of model size from network topology effects.